### v1

# Shock waves in nonlinear transmission lines

Abstract
We consider interaction between the small amplitude travelling waves (”sound”) and the shock waves in the transmission line containing both nonlinear capacitors and nonlinear inductors. We calculate for the ”sound” wave the coefficient of reflection from (the coefficient of transmission through) the shock wave. These coefficients are expressed in terms of the wave speeds and the wave impedances. When only the capacitors or only the inductors are nonlinear, the coefficients are expressed in terms of the wave speeds only. We explicitly include into consideration of the shocks the dissipation, introducing ohmic resistors shunting the inductors and also in series with the capacitors. This allows us to describe the shocks as physical objects of finite width and study their profiles. In some particular cases the profiles were obtained in terms of elementary functions.
I Introduction
The nonlinear electrical transmission lines are of much interest both due to their applications, and as the laboratories to study nonlinear waves malomed2 . A very interesting particular type of signals which can propagate along such lines - the shock waves - is attracting interest since long ago landauer1 ; landauer2 . We published a series of papers on the travelling waves in Josephson transmission lines (JTL): the kinks, the solitons and the shocks kogan4 (see the last publication of that series and the references therein). The cited publication is especially relevant for what will be presented below. In that publication we considered among other problems, the small amplitude wave (”sound”) reflection from and transmission through the shock wave.
Note that from the point of view of an electrical engineer, Josephson junctions are just the nonlinear inductors, and this is how they were described in all our papers on JTL. On the other hand, the capacitors in all these papers were assumed to be linear. On the other hand, very recently we published the paper on the travelling waves in the transmission line with the linear inductors but with the nonlinear capacitors kogan5 .
In the present paper we consider the travelling waves in the transmission line where, in the general case, both the inductors and the capacitors are nonlinear. In the first part of the paper, generalizing our previous results kogan4 ; kogan5 , we calculate the ”sound” reflection from and transmission through the shock wave in these nonlinear transmission lines. In the second part of the paper we include explicitly the ohmic losses (which are of course necessary for the existence of the shock waves) into the circuit equations. This allows us to study the profile of the shock wave.
II The circuit equations
The transmission line constructed from the identical nonlinear inductors and the identical nonlinear capacitors is shown on Fig. 1.
Figure 1: Lossless nonlinear transmission line
We take the voltages on the capacitors $V_{n}$ and the current through the inductors $I_{n}$ as the dynamical variables. The circuit equations (Kirchhoff laws) are
$\frac{dQ_{n}}{dt}$ $=I_{n}-I_{n+1},$ (1a)
$\frac{d\Phi_{n+1}}{dt}$ $=V_{n}-V_{n+1},$ (1b)
where
[…]
$Q_{n}$ is the charge of the nonlinear capacitor and ${\cal V}_{n}$ is the voltage on the nonlinear inductor.
In the continuum approximation we treat $n$ as the continuous variable $x$ and approximate the finite differences in the r.h.s. of the equations by the first derivatives with respect to $x$, after which the equations take the form
$\frac{\partial Q}{\partial t}$ $=-\frac{\partial I}{\partial x},$ (3a)
$\frac{\partial\Phi}{\partial t}$ $=-\frac{\partial V}{\partial x}$ (3b)
(we measure distance in the units of the transmission line period). To close the system (3) we should specify
$Q$ $=Q(V),$ (4a)
$\Phi$ $=\Phi(I).$ (4b)
Further on we’ll consider $Q$ and $\Phi$ as known functions of $V$ and $I$ respectively. Note however, that sometimes it will be convenient for us to invert the relations (4) and consider $V$ and $I$ as known functions of $Q$ and $\Phi$ respectively. We’ll switch between the representations without further notice.
III The ”sound” and the shocks
Because (3) is nonlinear, it can have a lot of different types of solutions. In this paper we’ll be interested in only two types of those. First type - small amplitude waves on a homogeneous background $\varphi_{0},V_{0}$.
$V$ $=V_{0}+v,$ (5a)
$I$ $=I_{0}+i.$ (5b)
For such waves Eq. (3) is simplified to
$C_{0}\frac{\partial v}{\partial t}$ $=-\frac{\partial i}{\partial x},$ (6a)
$L_{0}\frac{\partial i}{\partial t}$ $=-\frac{\partial v}{\partial x},$ (6b)
where
$C_{0}$ $=\left.\frac{dQ}{dV}\right|_{V=V_{0}},$ (7a)
$L_{0}$ $=\left.\frac{d\Phi}{dI}\right|_{I=I_{0}}.$ (7b)
The solutions of Eq. (6) (which, for brevity we’ll call the ”sound”) are right- and left-propagating travelling waves, each depending upon the single variable $\tau_{\pm}=t\mp x/u_{0}$, where
$u_{0}^{2}=\frac{1}{L_{0}C_{0}}$ (8)
is the ”sound” speed. The current and the voltage in the waves are connected by the equations
$v=\pm Z_{0}i,$ (9)
where
$Z_{0}=\sqrt{\frac{L_{0}}{C_{0}}}$ (10)
is the linear impedance of the transmission line.
The second type of solutions, we are interested in, is shock waves, satisfying the conditions (see Section V)
$U_{21}\left(Q_{2}-Q_{1}\right)$ $=I_{2}-I_{1},$ (11a)
$U_{21}\left(\Phi_{2}-\Phi_{1}\right)$ $=V_{2}-V_{1},$ (11b)
where the indices 1 and 2 refer to the quantities before and after the shock respectively, and $U_{21}$ is the shock wave speed. Equations (11) mean, in particular, that from the five parameters ($V_{1},V_{2},I_{1},I_{2},U$) - three are the free parameters, and the other two are defined by the equations.
The indices 1 and 2 enter in a symmetrical way in the system 11, so Eqs. 11 as they stand can describe propagation of the shock in both directions. However, in Section V) we’ll show that the shock always propagates in such a way that whitham ; kogan4 ; kogan5
$u_{2}^{2}>U_{21}^{2}>u_{1}^{2}.$ (12)
IV The ”sound” reflection and transmission
IV.1 Nonlinear transmission line
We’ll be interested in two problems landau ; congy . The first one: A ”sound” wave is incident from the rear on a shock wave (see Fig. 2).
Figure 2: Reflection of a ”sound” wave from a shock wave.
The second problem: A ”sound” wave is incident from the front on a shock wave (see Fig. 3).
Figure 3: Transmission of a ”sound” wave through a shock wave.
(Note the difference between the problems we consider and the problem of propagation of high-frequency wave packets along a smooth evolving background flow kamchatnov .) While formulating both problems we took into account the inequalities (12), meaning that there is no ”sound” transmission in the first case, and no ”sound” reflection in the second case.
For the first problem mentioned above we have
$\left(\begin{array}[]{c|c}V_{1}&V_{2}\\ I_{1}&I_{2}\end{array}\right)=\left(\begin{array}[]{c|c}V_{b}&V_{a}+v^{(in)}+v^{(r)}\\ I_{b}&I_{a}+i^{(in)}+i^{(r)}\end{array}\right),$ (17)
where (in) stands for the incident ”sound” wave and (r) for the reflected ”sound” wave. We also have to take into account the change of the shock speed due to the ”sound” waves
$U_{21}=U_{ab}+\delta U^{(refl)}.$ (18)
Substituting (17) into (11), in linear with respect to the ”sound” wave amplitudes approximation we obtain
$\delta U^{(refl)}\left(Q_{a}-Q_{b}\right)+U_{ab}C_{a}\left(v^{(in)}+v^{(r)}\right)=i^{(in)}+i^{(r)},$ (19a)
$\delta U^{(refl)}\left(\Phi_{a}-\Phi_{b}\right)+U_{ab}L_{a}\left(i^{(in)}+i^{(r)}\right)=v^{(in)}+v^{(r)}.$ (19b)
Taking into account the relations
$v^{(in)}$ $=Z_{a}i^{(in)},$ (20a)
$v^{(r)}$ $=-Z_{a}i^{(r)}$ (20b)
(the difference in the signs is because of the opposite directions of propagation of the two waves) and dividing the equations (19) to each other we obtain
$Z_{ab}=Z_{a}\frac{u_{in}i^{(in)}+\left(u_{a}+U_{ab}\right)i^{(r)}}{u_{r}i^{(in)}-\left(u_{a}+U_{ab}\right)i^{(r)}},$ (21)
where $u_{in}=u_{a}-U$ is the speed of the incident ”sound” wave relative to the shock wave, $u_{r}=u_{a}+U$ is the speed of the reflected ”sound” wave relative to the shock wave and we introduced the impedance of the shock wave
$Z_{ab}\equiv\frac{V_{a}-V_{b}}{I_{a}-I_{b}}=\frac{\Phi_{a}-\Phi_{b}}{Q_{a}-Q_{b}}.$ (22)
Solving (21) we obtain for the reflection coefficient
$R\equiv\frac{i^{(r)}}{i^{(in)}}=\frac{u_{in}\left(Z_{ab}-Z_{a}\right)}{u_{r}\left(Z_{ab}+Z_{a}\right)}.$ (23)
As one could have expected, the modulus of the ”sound” reflection coefficient is less than one, and it goes to zero when the intensity of the shock wave decreases, that is when $\Phi_{a}\to\Phi_{b},V_{a}\to V_{b}$, in other words, when the shock wave itself nearly becomes the ”sound” wave.
For the second problem we have
$\left(\begin{array}[]{c|c}V_{1}&V_{2}\\ I_{1}&I_{2}\end{array}\right)=\left(\begin{array}[]{c|c}V_{b}+V^{(in)}&V_{a}+v^{(t)}\\ I_{b}+i^{(in)}&I_{a}+i^{(t)}\end{array}\right),$ (28)
where (t) stands for the transmitted wave. Substituting (28) into (11), in linear approximation we obtain
$\delta U^{(trans)}\left(Q_{a}-Q_{b}\right)+U_{ab}\left(C_{a}v^{(t)}-C_{b}v^{(in)}\right)=i^{(t)}-i^{(in)},$ (29a)
$\delta U^{(trans)}\left(\Phi_{a}-\Phi_{b}\right)+U_{ab}\left(L_{a}i^{(t)}-L_{b}i^{(in)}\right)=v^{(t)}-v^{(in)}.$ (29b)
Taking into account the relations
$v^{(in)}$ $=-Z_{b}i^{(in)},$ (30a)
$v^{(t)}$ $=-Z_{a}i^{(t)}$ (30b)
and dividing the equations (29) to each other we obtain
$Z_{ab}=\frac{u_{a}u_{in}Z_{b}i^{(in)}-u_{b}u_{t}Z_{a}i^{(t)}}{-u_{a}u_{in}i^{(in)}+u_{b}u_{t}i^{(t)}},$ (31)
where $u_{in}=u_{b}+U$ is the speed of the incident ”sound” wave relative to the shock wave, and $u_{t}=u_{a}+U$ is the speed of the transmitted ”sound” wave relative to the shock wave. Solving (31) we obtain for the transmission coefficient
$T\equiv\frac{i^{(t)}}{i^{(in)}}=\frac{u_{a}u_{in}\left(Z_{ab}+Z_{b}\right)}{u_{b}u_{t}\left(Z_{ab}+Z_{a}\right)}.$ (32)
As one could have expected, the transmission coefficient goes to 1 when the intensity of the shock wave decreases, that is when $\Phi_{a}\to\Phi_{b},V_{a}\to V_{b}$, in other words, when the shock wave itself nearly becomes the ”sound” wave.
Note that Eqs. (21) and (31) can be written down respectively in the form
$u_{in}\left(Z_{ab}-Z_{a}\right)i^{(in)}$ $=u_{r}\left(Z_{ab}+Z_{a}\right)i^{(r)},$ (33a)
$u_{a}u_{in}\left(Z_{ab}+Z_{b}\right)i^{(in)}$ $=u_{b}u_{t}\left(Z_{ab}+Z_{a}\right)i^{(t)}.$ (33b)
We wonder whether these equations could have been written down at an earlier stage by considering propagation of the waves in the frame of reference moving with the shock wave.
IV.2 Half-nonlinear transmission line
IV.2.1 Linear capacitor
For linear capacitor ($Q=VC$, where $C$ is a constant) from (8) and (10) follows that
$Z_{a,b}=\frac{1}{Cu_{a,b}},$ (34)
and from (11) and (22) follows that
$Z_{ab}=\frac{1}{CU_{ab}}.$ (35)
Hence Eqs. (23) and (32) can be simplified respectively to
$R$ $=\frac{u_{in}^{2}}{u_{r}^{2}},$ (36a)
$T$ $=\frac{u_{a}^{2}u_{in}^{2}}{u_{b}^{2}u_{t}^{2}},$ (36b)
and from (8) and (11) follows
$u_{a,b}^{2}$ $=\frac{1}{C}\left.\frac{dI}{d\Phi}\right|_{\Phi=\Phi_{a,b}}$ (37a)
$U_{ab}^{2}$ $=\frac{1}{C}\frac{I_{a}-I_{b}}{\Phi_{a}-\Phi_{b}}.$ (37b)
Consider the case of JTL kogan4 . In this case from the second Josephson law
${\cal V}=\frac{\hbar}{2e}\frac{d\varphi}{dt},$ (38)
where $\varphi$ is Josephson phase, we obtain
$\Phi=\frac{\hbar}{2e}\varphi.$ (39)
Hence the first Josephson law gives
$I=I_{c}\sin\left(\frac{2e\Phi}{\hbar}\right),$ (40)
and we can easily calculate all the velocities. Equations (36) for the case of JTL were obtained in our previous publication kogan4 . (Comparing the results of Ref. kogan4 with those presented in this paper, one should pay attention to a slightly different definitions of the transmission and the reflection coefficients in the two papers.)
IV.2.2 Linear inductor
For linear inductor ($\Phi=LI$, where $L=$ is a constant) from (8) and (10) follows that
$Z_{a,b}=Lu_{a,b},$ (41)
and from (11) and (22) follows that
$Z_{ab}=LU_{ab}.$ (42)
Hence Eqs. (23) and (32) can be respectively simplified to
$R$ $=-\frac{u_{in}^{2}}{u_{r}^{2}},$ (43a)
$T$ $=\frac{u_{a}u_{in}^{2}}{u_{b}u_{t}^{2}},$ (43b)
and from (8) and (11) follows
$u_{a,b}^{2}$ $=\frac{1}{L}\left.\frac{dV}{dQ}\right|_{Q=Q_{a,b}}$ (44a)
$U_{ab}^{2}$ $=\frac{1}{L}\frac{V_{a}-V_{b}}{Q_{a}-Q_{b}}.$ (44b)
Note that because of the symmetry of Eq. (3) with respect to simultaneous interchange
$Q$ $\leftrightarrows\Phi,$ (45a)
$V$ $\leftrightarrows I,$ (45b)
Eqs. (43), (44) follow from Eqs. (36), (37).
V Lossy transmission line
V.1 Nonlinear transmission line
As it is well known landau , the shock waves exist only in the presence of dissipation. (In the absence of the dissipation, the localized travelling waves in the systems we consider are kinks and solitons kogan1 ; kogan2 ; kogan4 .) On the other hand, many properties of the shocks in fluids were studied beginning from the XIX century without taking the dissipation explicitly into account, but just postulating the existence of discontinuous solutions and the jump conditions krehl . This is exactly what we did in the previous Sections.
However in this Section, to be consistent, we assume the existence of ohmic resistors in the transmission line, shunting the inductors and in series with the capacitors. Such transmission line is shown on Fig. 4. We’ll resolve the zero width discontinuities from the previous Sections, which will turn out to be narrow (for weak dissipation) crossover regions and rigorously prove the jump conditions.
Figure 4: Lossy nonlinear transmission line
The corrected Eqs. (1) are
$\frac{dQ_{n}}{dt}$ $=I_{n}-I_{n+1}+\frac{1}{R_{L}}\frac{d}{dt}\left(\Phi_{n}-\Phi_{n+1}\right),$ (46a)
$\frac{d\Phi_{n+1}}{dt}$ $=V_{n}-V_{n+1}+R_{C}\frac{d}{dt}\left(Q_{n}-Q_{n+1}\right).$ (46b)
In the continuum approximation (46) takes the form (compare with (3))
$\frac{\partial Q}{\partial t}$ $=-\frac{\partial I}{\partial x}-\frac{1}{R_{L}}\frac{\partial^{2}\Phi}{\partial x\partial t},$ (47a)
$\frac{\partial\Phi}{\partial t}$ $=-\frac{\partial V}{\partial x}-R_{C}\frac{\partial^{2}Q}{\partial x\partial t}.$ (47b)
Further on we’ll consider the travelling waves, for which all the dependent variables depend upon the single independent variable $\tau=t-x/U$, where $U$ is the speed of the wave. For such waves (47) turns into the system of ODE which after integration takes the form
$\frac{1}{R_{L}}\frac{d\Phi}{d\tau}$ $=UQ-I(\Phi)+\text{const},$ (48a)
$R_{C}\frac{dQ}{d\tau}$ $=U\Phi-V(Q)+\text{const},$ (48b)
To study the shock waves we should impose the boundary conditions
$\lim_{\tau\to-\infty}Q(\tau)$ $=Q_{1};\hskip 14.22636pt\lim_{\tau\to+\infty}Q(\tau)=Q_{2};$ (49a)
$\lim_{\tau\to-\infty}\Phi(\tau)$ $=\Phi_{1};\hskip 14.22636pt\lim_{\tau\to+\infty}\Phi(\tau)=\Phi_{2}.$ (49b)
Both $(Q_{1},\Phi_{1})$ and $(Q_{2},\Phi_{2})$ are the fixed points (in the $(Q,\Phi)$ phase space), hence the r.h.s. of (48b) is equal to zero in both points. Equating these two values we recover the jump conditions (11b). Doing the same procedure with (48a) we recover the jump conditions (11a).
The point $(Q_{1},\Phi_{1})$ is the unstable fixed point of the system (48). This means
$\left|\begin{array}[]{cc}\frac{d}{d\Phi}\left[U_{21}Q-I(\Phi)\right]_{Q_{1},\Phi_{1}}&\frac{d}{dQ}\left[U_{21}Q-I(\Phi)\right]_{Q_{1},\Phi_{1}}\\ \frac{d}{d\Phi}\left[U_{21}\Phi-V(Q)\right]_{Q_{1},\Phi_{1}}&\frac{d}{dQ}\left[U_{21}\Phi-V(Q)\right]_{Q_{1},\Phi_{1}}\end{array}\right|<0,$ (52)
or
$\left|\begin{array}[]{cc}-\left.\frac{dI}{d\Phi}\right|_{\Phi_{1}}&U_{21}\\ U_{21}&-\left.\frac{dV}{dQ}\right|_{Q_{1}}\end{array}\right|<0.$ (56)
Calculating the determinant and recalling the result for the ”sound” speed we obtain
$U_{21}^{2}>u_{1}^{2}.$ (57)
On the other hand, the point $(Q_{2},\Phi_{2})$ is the stable fixed point, hence we obtain
$u_{2}^{2}>U_{21}^{2}.$ (58)
Thus we proved the inequalities (12) postulated previously.
V.2 Half-nonlinear transmission line
Further analysis of the fully nonlinear line we relegate to Appendix A) and here consider transmission line with the linear capacitor. Because of the symmetry of Eq. (48) given by (45) (to which we should add $1/R_{L}\leftrightarrows R_{C}$), the case of the transmission line with linear capacitors and nonlinear inductors (JTL first and foremost) would be identical to the one considered in this Section..
Because of the relation $\Phi=LI$, where $L$ is a constant, we can exclude the current from (48) and after a bit of algebra obtain
$\frac{LR_{C}}{R_{L}}\frac{d^{2}Q}{d\tau^{2}}+R_{C}\frac{dQ}{d\tau}+\frac{L}{R_{L}}\frac{dV}{d\tau}=F,$ (59)
where
$F\equiv\frac{V_{2}-V_{1}}{Q_{2}-Q_{1}}Q-V+\frac{V_{1}Q_{2}-V_{2}Q_{1}}{Q_{2}-Q_{1}}.$ (60)
Equation (59) describes motion with friction of a fictitious Newtonian particle, $Q$ being the coordinate of the particle, in the potential well defined by (60). The boundary conditions (49a) tell that the particle starts its motion at $\tau=-\infty$ at the potential maximum at $Q=Q_{1}$ and ends - at $\tau=+\infty$ at the potential minimum at $Q=Q_{2}$, thus it is crucial to keep the nonlinear dependence of $Q$ upon $V$ in 60. We may present these maximum/minimum conditions as
$\left.\frac{dF(Q)}{dQ}\right|_{Q=Q_{2}}=\frac{V_{2}-V_{1}}{Q_{2}-Q_{1}}-\left.\frac{dV}{dQ}\right|_{Q=Q_{2}}<0$ (61)
and opposite inequality for $V=V_{1}$. Recalling (44) we realise that we have reproduced in this particular case Eqs. (57), (58).
Equation (59) is simplified (in a trivial way) when there is a leading mechanism of dissipation, that is when either $R_{C}Q\ll LV/R_{L}$ or $R_{C}Q\gg LV/R_{L}$. In each of these cases, in the l.h.s. of (59) a single term with the first derivative should be left, and the equation can be easily integrated in quadratures (in some cases in elementary functions). Thus we obtained Eqs. (48a) and (48b) of our paper kogan3 .
V.2.1 Weak shocks
Alternatively, Eq. (59) is simplified (in a less trivial way) for weak shock, when $|V_{1}-V_{2}|$ is small. We can put in the l.h.s. of the equation
$\frac{dQ}{d\tau}=C_{s}\frac{dV}{d\tau},$ (62)
where $C_{s}$ is a differential capacitance assumed to be constant for the voltages considered, and (59) becomes
$T^{2}\frac{d^{2}V}{d\tau^{2}}+\tau_{R}\frac{dV}{d\tau}=F(V),$ (63)
where $\tau_{R}\equiv R_{C}C_{s}+L/R_{L}$ and $T^{2}\equiv LC_{s}R_{C}/R_{L}$. The solution of (63) should satisfy the boundary conditions (49a), rewritten as
$\lim_{\tau\to-\infty}V(\tau)=V_{1};\hskip 14.22636pt\lim_{\tau\to+\infty}V(\tau)=V_{2}.$ (64)
Before going further it is convenient to rewrite (60) as
$F(V)=\frac{(V-V_{1})\left[Q(V)-Q_{2}\right]-\left\{2\leftrightarrows 1\right\}}{Q_{2}-Q_{1}},$ (65)
where the notation $\left\{2\leftrightarrows 1\right\}$ means the preceding term with the indexes $2$ and $1$ interchanged. For weak shocks we can approximate $Q(V)$ in (65) as
$Q(V)=Q_{0}+C_{0}(V+\beta V^{2}),$ (66)
and (63) becomes
$T^{2}\frac{d^{2}V}{d\tau^{2}}+\tau_{R}\frac{dV}{d\tau}=\beta(V-V_{1})(V-V_{2}).$ (67)
In the Appendix B we’ll show that when the parameters of the problem satisfy the condition
$\frac{\tau_{R}^{2}}{T^{2}}=\frac{25\beta(V_{1}-V_{2})}{6},$ (68)
there exists the elementary solution of (67)
$V=V_{2}+\frac{V_{1}-V_{2}}{\left[\exp\left(\chi\tau\right)+1\right]^{2}},$ (69)
where
$\chi=\frac{\tau_{R}}{5T^{2}}.$ (70)
Approximation (66) can be inadequate, if $V_{1},V_{2}$ are close to zero and, as it normally happens, the charge is an odd function of the voltage. So we are motivated to generalize (66) to
$Q=Q_{0}+C_{0}(V+\beta V^{2}+\gamma V^{3}).$ (71)
In this case (63) becomes
$T^{2}\frac{d^{2}V}{d\tau^{2}}+\tau_{R}\frac{dV}{d\tau}=\gamma(V-V_{1})(V-V_{2})(V+V_{3}),$ (72)
where $V_{3}=\beta/\gamma+V_{1}+V_{2}$.
In the Appendix B we’ll show that when the parameters of the problem satisfy the condition
$\frac{\tau_{R}^{2}}{T^{2}}=\frac{\gamma\left(V_{1}+V_{2}+2V_{3}\right)^{2}}{2},$ (73)
there exists the elementary solution of (72)
$V=V_{2}+\frac{V_{1}-V_{2}}{\exp\left(\psi\tau\right)+1},$ (74)
where
$\psi=\frac{\tau_{R}\left(V_{1}-V_{2}\right)}{T^{2}\left(V_{1}+V_{2}+2V_{3}\right)}.$ (75)
Acknowledgements.
We are grateful to A. M. Kamchatnov for sending us the manuscript of his (to be published) excellent book ”Theory of Nonlinear Waves”, which motivated the continuation of our previous studies. We are very grateful to J. Cuevas-Maraver, M. Goldstein, A. Kamchatnov, B. Malomed, A. Novic-Cohen and V. Verbuk for the insightful discussions. We are also grateful to the organizers of Nor-Amberd School in Theoretical Physics, the scientific atmosphere of which stimulated the presented research. The paper was finalized during the Author’s participation in School and Workshop on Dynamical Systems, ICTP, Trieste. The Author is grateful to the Center for the hospitality.
Appendix A Shock profile for the fully nonlinear transmission line
Taking into account the boundary conditions we can present (48) as
$\frac{1}{R_{L}}\frac{d\Phi}{d\tau}$ $=UQ-I(\Phi)-UQ_{1}+I_{1},$ (76a)
$R_{C}\frac{dQ}{d\tau}$ $=U\Phi-V(Q)-U\Phi_{1}+V_{1},$ (76b)
Equation (76) allows us to study the profile of the shock wave in the general case.
Equation (76) is simplified when there is a leading mechanism of dissipation, in which cases we should substitute into (76) either $R_{C}=0$ or $R_{L}=\infty$. In the first case we equate the r.h.s. of (76b) to zero and obtain
$V(Q)=U\Phi-U\Phi_{1}+V_{1}.$ (77)
Solving (77) with respect to $Q$ and substituting into (76a) we obtain the equation which can be easily integrated in quadratures.
In the second case we equate the r.h.s. of (76a) to zero and obtain
$I(\Phi)=UQ-UQ_{1}+I_{1}.$ (78)
Solving (78) with respect to $\Phi$ and substituting into (76b) we obtain the equation which can be integrated in quadratures.
If both $R_{C}>0$ and $R_{L}<\infty$, from (76) we may obtain the equation
$\frac{1}{R_{L}R_{C}}\frac{d\Phi}{dQ}=\frac{UQ-I(\Phi)-UQ_{1}+I_{1}}{U\Phi-V(Q)-U\Phi_{1}+V_{1}}.$ (79)
If we can solve (79), we can substitute the solution into one of the equations (76) and integrate the resulting equations in quadratures to obtain the profile of the shock wave.
Appendix B Integration of the ODE which don’t contain explicitly the independent variable in terms of elementary functions
Looking at Eq. (63) we realize that it doesn’t contain explicitly the independent variable $\tau$. This prompts the idea to consider $V$ as the new independent variable and $E\equiv dV/d\tau$ as the new dependent variable. In the new variables (63) takes the form
$T^{2}E\frac{dE}{dV}+\tau_{R}E=F(V).$ (80)
As it follows from (49a), the boundary conditions for (80) are
$\lim_{V\to V_{1},V_{2}}E(V)=0.$ (81)
Note that Eq. (80) and the boundary conditions (81) are symmetric with respect to interchange of the indexes $2$ and $1$, while Eq. (63) and the boundary conditions (49a) - certainly not. So to the boundary conditions (64) we should add the demand (61) (or opposite inequality for $V=V_{1}$). Substituting (61) into (80) we obtain the demand written in the new variables as
$\left.\frac{T^{2}}{2}\frac{d^{2}E}{dV^{2}}\right|_{V=V_{2}}+\left.\tau_{R}\frac{dE}{dV}\right|_{V=V_{2}}<0.$ (82)
Since (80) is the first order differential equation, the problem is overdetermined. However, for the problems considered explicitly in Section V.2.1, not only will we be able to find the solutions of (80) satisfying the boundary conditions (81) - the solutions will turn out to be simple polynomial (fractional polynomial) functions.
For $Q(V)$ given by (71), Eq. (80) can be presented as
$T^{2}E\frac{dE}{dV}+\tau_{R}E=\gamma(V-V_{1})(V-V_{2})(V+V_{3}).$ (83)
The boundary conditions (81) and the r.h.s. of (83) prompt the idea to look for $E$ as the polynomial function proportional to $\left(V-V_{1}\right)\left(V-V_{2}\right)$. Because the r.h.s. of (83) is the third order polynomial of $V$, the multiplying factor should be a constant. That is the natural inspired guess is kogan4 ; kogan5
$E(V)=\frac{\psi}{V_{1}-V_{2}}\left(V-V_{1}\right)\left(V-V_{2}\right)$ (84)
(the normalization was chosen for later convenience).
Substituting (84) into (83) we obtain
$\frac{T^{2}\psi^{2}}{\left(V_{1}-V_{2}\right)^{2}}\left(2V-V_{1}-V_{2}\right)+\frac{\tau_{R}\psi}{V_{1}-V_{2}}=\gamma\left(V+V_{3}\right).$
Hence the condition on the parameters of the problem, for (84) to be the solution of (83) is given by (73), and the constant multiplier in (84) should is given by (75). Substituting (84) into the equation
$\frac{dV}{d\tau}=E(V)$ (86)
we obtain the solution (74).
For $Q(V)$ given by (66) we obtain, instead of (83), Abel equation of the second kind polyanin
$T^{2}E\frac{dE}{dV}+\tau_{R}E=\beta(V-V_{1})(V-V_{2}).$ (87)
Counting the powers of $V$ in the r.h.s. of (87) we come to the conclusion that if $E$ is a polynomial function of $V$, it should be a 3/2 order polynomial function. And the boundary conditions (81) tell us that this polynomial should zero out at $V=V_{1,2}$. To acquire an inspired guess how the polynomial may look like, let us return to the solution of (83) we obtained previously and split in two one multiplier in the solution
$\frac{V-V_{1}}{V_{1}-V_{2}}=\left[\left(\frac{V-V_{2}}{V_{1}-V_{2}}\right)^{1/2}-1\right]\left[\left(\frac{V-V_{2}}{V_{1}-V_{2}}\right)^{1/2}+1\right].$
All that was said above being taken together, prompts us to check up the solution of (87) in the form kogan4 ; kogan5
$E(V)=2\chi\left[\left(\frac{V-V_{2}}{V_{1}-V_{2}}\right)^{1/2}-1\right]\left(V-V_{2}\right).$ (89)
Substituting the ansatz into (87) we see that it is in fact the solution of the equation, provided the parameters of the problem satisfy the relation (68) and $\chi$ is given by (70). Substituting (89) into (86) we obtain (69)
Actually, because of the symmetry of Eq. (80) and the boundary conditions (81) with respect to interchange of the indexes $2$ and $1$, instead of (89) and the condition (68) we could as well have obtained those with the indexes $2$ and $1$ interchanged. However, one can easily see that (89) satisfies the demand (82), and (89) with the indexes 1 and 2 interchanged - doesn’t.

### v3

@@
 (we measure distance in the units of the transmission line period). To close the system (3) we should specify
-$Q$ $=Q(V),$ (4a)
-$\Phi$ $=\Phi(I).$ (4b)
-Further on we’ll consider $Q$ and $\Phi$ as known functions of $V$ and $I$ respectively.
-Note however, that sometimes it will be convenient for us to invert the relations (4) and consider $V$ and $I$ as known functions of $Q$ and $\Phi$ respectively.
-We’ll switch between the representations without further notice.
+$V$ $=V(Q),$ (4a)
+$I$ $=\Phi(\Phi).$ (4b)
+Further on we’ll consider $V$ and $I$ as known functions of $Q$ and $\Phi$ respectively.
 III The ”sound” and the shocks
@@
 where
-$C_{0}$ $=\left.\frac{dQ}{dV}\right|_{V=V_{0}},$ (7a)
-$L_{0}$ $=\left.\frac{d\Phi}{dI}\right|_{I=I_{0}}.$ (7b)
+$C_{0}$ $=\left.\frac{dQ}{dV}\right|_{Q=Q_{0}},$ (7a)
+$L_{0}$ $=\left.\frac{d\Phi}{dI}\right|_{\Phi=\Phi_{0}}.$ (7b)
 The solutions of Eq. (6) (which, for brevity we’ll call the ”sound”) are right- and left-propagating travelling waves, each depending upon the single variable $\tau_{\pm}=t\mp x/u_{0}$, where
@@
 is the linear impedance of the transmission line.
-The second type of solutions, we are interested in, is shock waves, satisfying the conditions (see Section V)
+The second type of solutions, we are interested in, is shock waves, containing (in the approximation of the present Section) discontinuities, satisfying the conditions (see Section V)
 $U_{21}\left(Q_{2}-Q_{1}\right)$ $=I_{2}-I_{1},$ (11a)
 $U_{21}\left(\Phi_{2}-\Phi_{1}\right)$ $=V_{2}-V_{1},$ (11b)
-where the indices 1 and 2 refer to the quantities before and after the shock respectively, and $U_{21}$ is the shock wave speed.
-Equations (11) mean, in particular, that from the five parameters ($V_{1},V_{2},I_{1},I_{2},U$) - three are the free parameters, and the other two are defined by the equations.
-The indices 1 and 2 enter in a symmetrical way in the system 11, so Eqs.
-11 as they stand can describe propagation of the shock in both directions.
-However, in Section V) we’ll show that the shock always propagates in such a way that whitham ; kogan4 ; kogan5
+where
+$V_{a,b}\equiv V(Q_{a,b}),\hskip 28.45274ptI_{a,b}\equiv I(\Phi_{a,b}),$ (12)
+and $U_{21}$ is the shock wave speed.
+The indices 1 and 2, which refer to the quantities before and after the shock respectively, enter in the system (11) in a symmetrical way.
+However, in Section V we’ll show that the shock always propagates in such a way that whitham ; kogan4 ; kogan5
 $u_{2}^{2}>U_{21}^{2}>u_{1}^{2}.$ (13)
+The shock is supersonic with respect to the phase before the front, but subsonic with respect to the phase behind the front.
 IV The ”sound” reflection and transmission
-IV.1 Nonlinear transmission line
 We’ll be interested in two problems landau ; congy . The first one: A ”sound” wave is incident from the rear on a shock wave (see Fig. 2).
@@
 $\left(\begin{array}[]{c|c}V_{1}&V_{2}\\ I_{1}&I_{2}\end{array}\right)=\left(\begin{array}[]{c|c}V_{b}&V_{a}+v^{(in)}+v^{(r)}\\ I_{b}&I_{a}+i^{(in)}+i^{(r)}\end{array}\right),$ (18)
-where (in) stands for the incident ”sound” wave and (r) for the reflected ”sound” wave. We also have to take into account the change of the shock speed due to the ”sound” waves
+where (in) stands for the incident and (r) - for the reflected ”sound” wave. We also have to take into account the change of the shock speed due to the ”sound” waves
 $U_{21}=U_{ab}+\delta U^{(refl)}.$ (19)
@@
 $\delta U^{(refl)}\left(\Phi_{a}-\Phi_{b}\right)+U_{ab}L_{a}\left(i^{(in)}+i^{(r)}\right)=v^{(in)}+v^{(r)}.$ (20b)
+Dividing the equations (20) to each other we obtain
+$u_{in}v^{(in)}-u_{r}v^{(r)}=Z_{ab}\left(u_{in}i^{(in)}-u_{r}i^{(r)}\right),$ (21)
+where $u_{in}=u_{a}-U_{ab}$ is the speed of the incident ”sound” wave relative to the shock wave, $u_{r}=u_{a}+U_{ab}$ is the speed of the reflected ”sound” wave relative to the shock wave and we introduced the impedance of the shock wave
+$Z_{ab}\equiv\frac{V_{a}-V_{b}}{I_{a}-I_{b}}=\frac{\Phi_{a}-\Phi_{b}}{Q_{a}-Q_{b}}.$ (22)
 Taking into account the relations
@@
 $v^{(r)}$ $=-Z_{a}i^{(r)}$ (23b)
-(the difference in the signs is because of the opposite directions of propagation of the two waves) and dividing the equations (19) to each other we obtain
-$Z_{ab}=Z_{a}\frac{u_{in}i^{(in)}+\left(u_{a}+U_{ab}\right)i^{(r)}}{u_{r}i^{(in)}-\left(u_{a}+U_{ab}\right)i^{(r)}},$ (21)
-where $u_{in}=u_{a}-U$ is the speed of the incident ”sound” wave relative to the shock wave, $u_{r}=u_{a}+U$ is the speed of the reflected ”sound” wave relative to the shock wave and we introduced the impedance of the shock wave
-$Z_{ab}\equiv\frac{V_{a}-V_{b}}{I_{a}-I_{b}}=\frac{\Phi_{a}-\Phi_{b}}{Q_{a}-Q_{b}}.$ (22)
-Solving (21) we obtain for the reflection coefficient
-$R\equiv\frac{i^{(r)}}{i^{(in)}}=\frac{u_{in}\left(Z_{ab}-Z_{a}\right)}{u_{r}\left(Z_{ab}+Z_{a}\right)}.$ (23)
-As one could have expected, the modulus of the ”sound” reflection coefficient is less than one, and it goes to zero when the intensity of the shock wave decreases, that is when $\Phi_{a}\to\Phi_{b},V_{a}\to V_{b}$, in other words, when the shock wave itself nearly becomes the ”sound” wave.
+(the difference in the signs is because of the opposite directions of propagation of the two waves) and solving (21) we obtain for the reflection coefficient
+$R\equiv\frac{i^{(r)}}{i^{(in)}}=\frac{\left(Z_{ab}-Z_{a}\right)u_{in}}{\left(Z_{ab}+Z_{a}\right)u_{r}}.$ (24)
 For the second problem we have
@@
 $\delta U^{(trans)}\left(\Phi_{a}-\Phi_{b}\right)+U_{ab}\left(L_{a}i^{(t)}-L_{b}i^{(in)}\right)=v^{(t)}-v^{(in)}.$ (30b)
+Dividing the equations (30) to each other we obtain
+$u_{a}u_{in}v^{(in)}-u_{b}u_{t}v^{(t)}=Z_{ab}\left(u_{a}u_{in}i^{(in)}-u_{b}u_{t}i^{(t)}\right),$
+where $u_{in}=u_{b}+U_{ab}$ is the speed of the incident ”sound” wave relative to the shock wave, and $u_{t}=u_{a}+U_{ab}$ is the speed of the transmitted ”sound” wave relative to the shock wave.
 Taking into account the relations
@@
 $v^{(t)}$ $=-Z_{a}i^{(t)}$ (32b)
-and dividing the equations (29) to each other we obtain
-$Z_{ab}=\frac{u_{a}u_{in}Z_{b}i^{(in)}-u_{b}u_{t}Z_{a}i^{(t)}}{-u_{a}u_{in}i^{(in)}+u_{b}u_{t}i^{(t)}},$ (31)
-where $u_{in}=u_{b}+U$ is the speed of the incident ”sound” wave relative to the shock wave, and $u_{t}=u_{a}+U$ is the speed of the transmitted ”sound” wave relative to the shock wave.
-Solving (31) we obtain for the transmission coefficient
-$T\equiv\frac{i^{(t)}}{i^{(in)}}=\frac{u_{a}u_{in}\left(Z_{ab}+Z_{b}\right)}{u_{b}u_{t}\left(Z_{ab}+Z_{a}\right)}.$ (32)
-As one could have expected, the transmission coefficient goes to 1 when the intensity of the shock wave decreases, that is when $\Phi_{a}\to\Phi_{b},V_{a}\to V_{b}$, in other words, when the shock wave itself nearly becomes the ”sound” wave.
-Note that Eqs. (21) and (31) can be written down respectively in the form
-$u_{in}\left(Z_{ab}-Z_{a}\right)i^{(in)}$ $=u_{r}\left(Z_{ab}+Z_{a}\right)i^{(r)},$ (33a)
-$u_{a}u_{in}\left(Z_{ab}+Z_{b}\right)i^{(in)}$ $=u_{b}u_{t}\left(Z_{ab}+Z_{a}\right)i^{(t)}.$ (33b)
-We wonder whether these equations could have been written down at an earlier stage by considering propagation of the waves in the frame of reference moving with the shock wave.
-IV.2 Half-nonlinear transmission line
-IV.2.1 Linear capacitor
+and solving (IV) we obtain for the transmission coefficient
+$T\equiv\frac{i^{(t)}}{i^{(in)}}=\frac{\left(Z_{ab}+Z_{b}\right)u_{a}u_{in}}{\left(Z_{ab}+Z_{a}\right)u_{b}u_{t}}.$ (33)
+As one could have expected, the modulus of the ”sound” reflection coefficient is less than one, and it goes to zero when the intensity of the shock wave decreases, that is when $Q_{a}\to Q_{b},\Phi_{a}\to\Phi_{b}$, in other words, when the shock wave itself is nearly the ”sound” wave.
+Similarly, in these circumstances the transmission coefficient goes to 1.
+Looking back at (21) and (IV) we wonder whether these equations could have been written down by considering propagation of the waves in the frame of reference moving with the shock wave.
+IV.1 Half-nonlinear transmission line
+IV.1.1 Linear capacitor
 For linear capacitor ($Q=VC$, where $C$ is a constant) from (8) and (10) follows that
@@
 kogan4 with those presented in this paper, one should pay attention to a slightly different definitions of the transmission and the reflection coefficients in the two papers.)
-IV.2.2 Linear inductor
+IV.1.2 Linear inductor
 For linear inductor ($\Phi=LI$, where $L=$ is a constant) from (8) and (10) follows that
@@
 V Lossy transmission line
-V.1 Nonlinear transmission line
 As it is well known landau , the shock waves exist only in the presence of dissipation.
@@
 $\lim_{\tau\to-\infty}\Phi(\tau)$ $=\Phi_{1};\hskip 14.22636pt\lim_{\tau\to+\infty}\Phi(\tau)=\Phi_{2}.$ (49b)
+From the five parameters ($Q_{1},Q_{2},\Phi_{1},\Phi_{2},U$) - three are arbitrary constants, and the other two are defined by the equations.
 Both $(Q_{1},\Phi_{1})$ and $(Q_{2},\Phi_{2})$ are the fixed points (in the $(Q,\Phi)$ phase space), hence the r.h.s.
@@
 Doing the same procedure with (48a) we recover the jump conditions (11a).
-The point $(Q_{1},\Phi_{1})$ is the unstable fixed point of the system (48). This means
+$(Q_{1},\Phi_{1})$ is the semistable fixed point of the system (48), and $(Q_{2},\Phi_{2})$ is the stable fixed point. (Other options would have demanded fine tuning of the parameters.) This means
 $\left|\begin{array}[]{cc}\frac{d}{d\Phi}\left[U_{21}Q-I(\Phi)\right]_{Q_{1},\Phi_{1}}&\frac{d}{dQ}\left[U_{21}Q-I(\Phi)\right]_{Q_{1},\Phi_{1}}\\ \frac{d}{d\Phi}\left[U_{21}\Phi-V(Q)\right]_{Q_{1},\Phi_{1}}&\frac{d}{dQ}\left[U_{21}\Phi-V(Q)\right]_{Q_{1},\Phi_{1}}\end{array}\right|<0,$ (52)
@@
 $U_{21}^{2}>u_{1}^{2}.$ (57)
-On the other hand, the point $(Q_{2},\Phi_{2})$ is the stable fixed point, hence we obtain
+Similarly for the point $(Q_{2},\Phi_{2})$ we obtain
 $u_{2}^{2}>U_{21}^{2}.$ (58)
 Thus we proved the inequalities (13) postulated previously.
-V.2 Half-nonlinear transmission line
-Further analysis of the fully nonlinear line we relegate to Appendix A) and here consider transmission line with the linear capacitor.
+Note that the kinks, which we considered in our previous publications in addition to shocks but do not consider here, correspond to both fixed points being semistable.
+Thus the kinks turn out to be supersonic both with respect to the phase before the front and the phase behind the front.
+V.1 Half-nonlinear transmission line
+Further analysis of the fully nonlinear line we relegate to Appendix B) and here consider transmission line with the linear capacitor.
 Because of the symmetry of Eq.
@@
 Because of the relation $\Phi=LI$, where $L$ is a constant, we can exclude the current from (48) and after a bit of algebra obtain
-$\frac{LR_{C}}{R_{L}}\frac{d^{2}Q}{d\tau^{2}}+R_{C}\frac{dQ}{d\tau}+\frac{L}{R_{L}}\frac{dV}{d\tau}=F,$ (59)
+$\frac{LR_{C}}{R_{L}}\frac{d^{2}Q}{d\tau^{2}}+R_{C}\frac{dQ}{d\tau}+\frac{L}{R_{L}}\frac{dV(Q)}{d\tau}=F(Q),$ (59)
 where
-$F\equiv\frac{V_{2}-V_{1}}{Q_{2}-Q_{1}}Q-V+\frac{V_{1}Q_{2}-V_{2}Q_{1}}{Q_{2}-Q_{1}}.$ (60)
+$F(Q)\equiv\frac{V_{2}-V_{1}}{Q_{2}-Q_{1}}Q-V(Q)+\frac{V_{1}Q_{2}-V_{2}Q_{1}}{Q_{2}-Q_{1}}.$ (60)
+To (59) we should add the boundary conditions (49a). (The third arbitrary constant appears when we connect $\Phi$ with the found $Q$.)
 Equation (59) describes motion with friction of a fictitious Newtonian particle, $Q$ being the coordinate of the particle, in the potential well defined by (60).
-The boundary conditions (49a) tell that the particle starts its motion at $\tau=-\infty$ at the potential maximum at $Q=Q_{1}$ and ends - at $\tau=+\infty$ at the potential minimum at $Q=Q_{2}$, thus it is crucial to keep the nonlinear dependence of $Q$ upon $V$ in 60.
+The boundary conditions (49a) tell that the particle starts its motion at $\tau=-\infty$ at the potential maximum at $Q=Q_{1}$ and ends - at $\tau=+\infty$ at the potential minimum at $Q=Q_{2}$.
 We may present these maximum/minimum conditions as
@@
 and opposite inequality for $V=V_{1}$. Recalling (44) we realise that we have reproduced in this particular case Eqs. (57), (58).
-Equation (59) is simplified (in a trivial way) when there is a leading mechanism of dissipation, that is when either $R_{C}Q\ll LV/R_{L}$ or $R_{C}Q\gg LV/R_{L}$.
-In each of these cases, in the l.h.s.
-of (59) a single term with the first derivative should be left, and the equation can be easily integrated in quadratures (in some cases in elementary functions).
-Thus we obtained Eqs.
-(48a) and (48b) of our paper kogan3 .
-V.2.1 Weak shocks
-Alternatively, Eq. (59) is simplified (in a less trivial way) for weak shock, when $|V_{1}-V_{2}|$ is small. We can put in the l.h.s. of the equation
-$\frac{dQ}{d\tau}=C_{s}\frac{dV}{d\tau},$ (62)
-where $C_{s}$ is a differential capacitance assumed to be constant for the voltages considered, and (59) becomes
-$T^{2}\frac{d^{2}V}{d\tau^{2}}+\tau_{R}\frac{dV}{d\tau}=F(V),$ (63)
-where $\tau_{R}\equiv R_{C}C_{s}+L/R_{L}$ and $T^{2}\equiv LC_{s}R_{C}/R_{L}$. The solution of (63) should satisfy the boundary conditions (49a), rewritten as
-$\lim_{\tau\to-\infty}V(\tau)=V_{1};\hskip 14.22636pt\lim_{\tau\to+\infty}V(\tau)=V_{2}.$ (64)
+V.1.1 Weak shocks
 Before going further it is convenient to rewrite (60) as
-$F(V)=\frac{(V-V_{1})\left[Q(V)-Q_{2}\right]-\left\{2\leftrightarrows 1\right\}}{Q_{2}-Q_{1}},$ (65)
-where the notation $\left\{2\leftrightarrows 1\right\}$ means the preceding term with the indexes $2$ and $1$ interchanged. For weak shocks we can approximate $Q(V)$ in (65) as
-$Q(V)=Q_{0}+C_{0}(V+\beta V^{2}),$ (66)
-and (63) becomes
-$T^{2}\frac{d^{2}V}{d\tau^{2}}+\tau_{R}\frac{dV}{d\tau}=\beta(V-V_{1})(V-V_{2}).$ (67)
-In the Appendix B we’ll show that when the parameters of the problem satisfy the condition
-$\frac{\tau_{R}^{2}}{T^{2}}=\frac{25\beta(V_{1}-V_{2})}{6},$ (68)
+$F(Q)=\frac{\left[V(Q)-V_{1}\right]\left(Q-Q_{2}\right)-\left\{2\leftrightarrows 1\right\}}{Q_{2}-Q_{1}},$ (62)
+where the notation $\left\{2\leftrightarrows 1\right\}$ means the preceding term with the indexes $2$ and $1$ interchanged.
+For weak shocks, when $|Q_{1}-Q_{2}|$ is small, we can approximate $V(Q)$ for $Q$ between $Q_{1}$ and $Q_{2}$ as
+$V(Q)=\text{const}+\frac{Q+\beta Q^{2}}{C_{s}},$ (63)
+where $\beta$ and $C_{s}$ are assumed to be constant. We also may take $V(Q)$ in the last term in the l.h.s. of (59) in the linear approximation. So finally Eq. (59) becomes
+$T^{2}\frac{d^{2}Q}{d\tau^{2}}+\tau_{R}\frac{dQ}{d\tau}=\beta(Q-Q_{1})(Q-Q_{2}),$ (64)
+where
+$\tau_{R}$ $\equiv R_{C}C_{s}+\frac{L}{R_{L}},$ (65a)
+$T^{2}$ $\equiv\frac{LC_{s}R_{C}}{R_{L}}.$ (65b)
+One can check up that when the parameters of the problem satisfy the condition
+$\frac{\tau_{R}^{2}}{T^{2}}=\frac{25\beta(Q_{1}-Q_{2})}{6},$ (66)
 there exists the elementary solution of (64)
-$V=V_{2}+\frac{V_{1}-V_{2}}{\left[\exp\left(\chi\tau\right)+1\right]^{2}},$ (69)
+$Q=Q_{2}+\frac{Q_{1}-Q_{2}}{\left[\exp\left(\chi\tau\right)+1\right]^{2}},$ (67)
 where
 $\chi=\frac{\tau_{R}}{5T^{2}}.$ (68)
-Approximation (66) can be inadequate, if $V_{1},V_{2}$ are close to zero and, as it normally happens, the charge is an odd function of the voltage. So we are motivated to generalize (66) to
-$Q=Q_{0}+C_{0}(V+\beta V^{2}+\gamma V^{3}).$ (71)
+Approximation (63) can be inadequate, if $Q_{1},Q_{2}$ are close to zero and, as it normally happens, the charge is an odd function of the voltage. So we are motivated to generalize (63) to
+$V(Q)=\text{const}+\frac{Q+\beta Q^{2}+\gamma Q^{3}}{C_{s}}.$ (69)
 In this case (59) becomes
-$T^{2}\frac{d^{2}V}{d\tau^{2}}+\tau_{R}\frac{dV}{d\tau}=\gamma(V-V_{1})(V-V_{2})(V+V_{3}),$ (72)
-where $V_{3}=\beta/\gamma+V_{1}+V_{2}$.
-In the Appendix B we’ll show that when the parameters of the problem satisfy the condition
-$\frac{\tau_{R}^{2}}{T^{2}}=\frac{\gamma\left(V_{1}+V_{2}+2V_{3}\right)^{2}}{2},$ (73)
+$T^{2}\frac{d^{2}Q}{d\tau^{2}}+\tau_{R}\frac{dQ}{d\tau}=\gamma(Q-Q_{1})(Q-Q_{2})(Q+Q_{3}),$ (70)
+where $Q_{3}=\beta/\gamma+Q_{1}+Q_{2}$. One can check up that when the parameters of the problem satisfy the condition
+$\frac{\tau_{R}^{2}}{T^{2}}=\frac{\gamma\left(Q_{1}+Q_{2}+2Q_{3}\right)^{2}}{2},$ (71)
 there exists the elementary solution of (70)
-$V=V_{2}+\frac{V_{1}-V_{2}}{\exp\left(\psi\tau\right)+1},$ (74)
+$Q=Q_{2}+\frac{Q_{1}-Q_{2}}{\exp\left(\psi\tau\right)+1},$ (72)
 where
-$\psi=\frac{\tau_{R}\left(V_{1}-V_{2}\right)}{T^{2}\left(V_{1}+V_{2}+2V_{3}\right)}.$ (75)
+$\psi=\frac{\tau_{R}\left(Q_{1}-Q_{2}\right)}{T^{2}\left(Q_{1}+Q_{2}+2Q_{3}\right)}.$ (73)
+In the Appendix A we’ll show how one arrives to the solutions (67) and (72).
 Acknowledgements.
@@
 Goldstein, A.
-Kamchatnov, B.
+Kamchatnov, E.
+Konig, B.
 Malomed, A.
@@
 The Author is grateful to the Center for the hospitality.
-Appendix A Shock profile for the fully nonlinear transmission line
-Taking into account the boundary conditions we can present (48) as
-$\frac{1}{R_{L}}\frac{d\Phi}{d\tau}$ $=UQ-I(\Phi)-UQ_{1}+I_{1},$ (76a)
-$R_{C}\frac{dQ}{d\tau}$ $=U\Phi-V(Q)-U\Phi_{1}+V_{1},$ (76b)
-Equation (76) allows us to study the profile of the shock wave in the general case.
+Appendix A The ODE which doesn’t contain explicitly the independent variable
+Equation (59) doesn’t contain explicitly the independent variable $\tau$. This prompts the idea to consider $Q$ as the new independent variable and
+${\cal I}=\frac{dQ}{d\tau}$ (74)
+as the new dependent variable. In the new variables (59) takes the form of Abel equation of the second kind polyanin
+$T^{2}{\cal I}\frac{d{\cal I}}{dQ}+\tau_{R}{\cal I}=F(Q).$ (75)
+As it follows from (49a), the boundary conditions for (75) are
+$\lim_{Q\to Q_{1},Q_{2}}{\cal I}(Q)=0.$ (76)
+Since (75) is the first order differential equation, the problem is overdetermined.
+However, for the problems considered explicitly in Section V.1.1, not only will we be able to find the solutions of (75) satisfying the boundary conditions (76) - the solutions will turn out to be simple polynomial (fractional polynomial) functions.
+A.1 Integration in terms of elementary functions
+For $Q(V)$ given by (69), Eq. (75) can be presented as
+$T^{2}{\cal I}\frac{d{\cal I}}{dQ}+\tau_{R}{\cal I}=\gamma(Q-Q_{1})(Q-Q_{2})(Q+Q_{3}).$ (77)
+The boundary conditions (76) and the r.h.s.
+of (77) prompt the idea kogan4 ; kogan5 to look for ${\cal I}(Q)$ as the polynomial function proportional to $\left(Q-Q_{1}\right)\left(Q-Q_{2}\right)$.
+Because the r.h.s.
+of (77) is the third order polynomial of $V$, the multiplying factor should be a constant
+${\cal I}(Q)=\frac{\psi}{Q_{1}-Q_{2}}\left(Q-Q_{1}\right)\left(Q-Q_{2}\right)$ (78)
+(the normalization was chosen for later convenience).
+Substituting (78) into (77) we obtain
+$\frac{T^{2}\psi^{2}}{\left(Q_{1}-Q_{2}\right)^{2}}\left(2Q-Q_{1}-Q_{2}\right)+\frac{\tau_{R}\psi}{Q_{1}-Q_{2}}=\gamma\left(Q+Q_{3}\right).$
+Hence the condition on the parameters of the problem, for (78) to be the solution of (77) is given by (71), and the constant multiplier in (78) should is given by (73).
+Substituting (78) into Eq.
+(74) we obtain the solution (72).
+For $Q(V)$ given by (63) we obtain, instead of (77),
+$T^{2}{\cal I}\frac{d{\cal I}}{dQ}+\tau_{R}{\cal I}=\beta(Q-Q_{1})(Q-Q_{2}).$ (80)
+If ${\cal I}$ is a polynomial function of $Q$, it should be a 3/2 order polynomial. And the boundary conditions (76) tell us that this polynomial should zero out at $Q=Q_{1,2}$.
+Let us split in two one multiplier in the solution (78)
+$\frac{Q-Q_{1}}{Q_{1}-Q_{2}}=\left[\left(\frac{Q-Q_{2}}{Q_{1}-Q_{2}}\right)^{1/2}-1\right]\left[\left(\frac{Q-Q_{2}}{Q_{1}-Q_{2}}\right)^{1/2}+1\right].$
+This prompts us the idea to look up for the solution of (80) in the form kogan4 ; kogan5
+${\cal I}(Q)=2\chi\left[\left(\frac{Q-Q_{2}}{Q_{1}-Q_{2}}\right)^{1/2}-1\right]\left(Q-Q_{2}\right).$ (82)
+Substituting the ansatz into (80) we see that it is in fact the solution of the equation, provided the parameters of the problem satisfy the relation (66) and $\chi$ is given by (68).
+Substituting thus obtained ${\cal I}(Q)$ into (74) we obtain (67)
+A.2 Solitary wave solution of a generalized KdV equation
+To additionally illustrate the method of integration used in this Appendix, let us recover in the framework of the method the solitary wave solution of a generalized KdV equation drazin
+$\phi_{t}+(n+1)(n+2)\phi^{n}\phi_{x}+\phi_{xxx}=0.$ (83)
+For a travelling wave solution
+$\phi(t,x)=\phi(\tau),\hskip 28.45274pt\tau=t-x/U,$ (84)
+the equation in partial derivatives (83) is reduced to the ODE
+$U^{3}\frac{d\phi}{d\tau}-U^{2}(n+1)(n+2)\phi^{n}\frac{d\phi}{d\tau}-\frac{d^{3}\phi}{d\tau^{3}}=0.$ (85)
+Equation (85) doesn’t contain explicitly the independent variable $\tau$. This prompts the idea to consider $\phi$ as the new independent variable and
+$E=\frac{d\phi}{d\tau}$ (86)
+as the new dependent variable. As the result, we can present the nonlinear third order ODE (85) as the linear second order ODE for $E^{2}$
+$\frac{d^{2}E^{2}}{d\phi^{2}}=2U^{3}-2(n+1)(n+2)U^{2}\phi^{n}.$ (87)
+A solitary wave corresponds to the boundary conditions
+$\lim_{\tau\to\pm\infty}\phi(\tau)=0,$ (88)
+which in the new variables are kogan5
+$E^{2}(0)=0,\hskip 28.45274pt\left.\frac{dE^{2}}{d\phi}\right|_{\phi=0}=0.$ (89)
+Solving (87) with these boundary conditions we obtain
+$E^{2}=U^{3}\phi^{2}-2U^{2}\phi^{n+2}.$ (90)
+Substituting thus obtained $E(\phi)$ into (86) and integrating we obtain the well known result drazin
+$\phi^{n}=\frac{U}{2}\operatorname{sech}^{2}\left[\frac{n\sqrt{U}}{2}\left(x-Ut\right)\right].$ (91)
+Appendix B Shock profile for the fully nonlinear transmission line
+For the fully nonlinear transmission line it is convenient to present the problem in the form slightly different from that used everywhere above.
+We’ll consider the shock speed $U$ and the conditions before the shock as given.
+Thus the boundary conditions are
+$\lim_{\tau\to-\infty}Q(\tau)=Q_{1};\hskip 14.22636pt\lim_{\tau\to-\infty}\Phi(\tau)$ $=\Phi_{1}$ (92)
+and we can concretize (48) to
+$\frac{1}{R_{L}}\frac{d\Phi}{d\tau}$ $=U(Q-Q_{1})-I+I(\Phi_{1}),$ (93a)
+$R_{C}\frac{dQ}{d\tau}$ $=U(\Phi-\Phi_{1})-V+V(Q_{1}).$ (93b)
 Equation (93) is simplified when there is a leading mechanism of dissipation, in which cases we should substitute into (93) either $R_{C}=0$ or $R_{L}=\infty$.
@@
 of (93b) to zero and obtain
-$V(Q)=U\Phi-U\Phi_{1}+V_{1}.$ (77)
-Solving (77) with respect to $Q$ and substituting into (76a) we obtain the equation which can be easily integrated in quadratures.
-In the second case we equate the r.h.s. of (76a) to zero and obtain
-$I(\Phi)=UQ-UQ_{1}+I_{1}.$ (78)
-Solving (78) with respect to $\Phi$ and substituting into (76b) we obtain the equation which can be integrated in quadratures.
+$V=U(\Phi-\Phi_{1})+V(Q_{1}).$ (94)
+Substituting $V$ into (4a), solving the equation with respect to $Q$ and then substituting thus obtained $Q(\Phi)$ into (93a) we obtain the equation which can be easily integrated in quadratures.
+In the second case we equate the r.h.s.
+of (93a) to zero and obtain
+$I=U(Q-Q_{1})+(\Phi_{1}).$ (95)
+Substituting $I$ into (4b), solving the equation with respect to $\Phi$ and then substituting thus obtained $\Phi(Q)$ into (93b) we obtain the equation which can be integrated in quadratures.
+Thus we obtained Eqs.
+(48a) and (48b) of Ref.
+kogan3 .
 If both $R_{C}>0$ and $R_{L}<\infty$, from (93) we may obtain the equation
-$\frac{1}{R_{L}R_{C}}\frac{d\Phi}{dQ}=\frac{UQ-I(\Phi)-UQ_{1}+I_{1}}{U\Phi-V(Q)-U\Phi_{1}+V_{1}}.$ (79)
-If we can solve (79), we can substitute the solution into one of the equations (76) and integrate the resulting equations in quadratures to obtain the profile of the shock wave.
-Appendix B Integration of the ODE which don’t contain explicitly the independent variable in terms of elementary functions
-Looking at Eq.
-(63) we realize that it doesn’t contain explicitly the independent variable $\tau$.
-This prompts the idea to consider $V$ as the new independent variable and $E\equiv dV/d\tau$ as the new dependent variable.
-In the new variables (63) takes the form
-$T^{2}E\frac{dE}{dV}+\tau_{R}E=F(V).$ (80)
-As it follows from (49a), the boundary conditions for (80) are
-$\lim_{V\to V_{1},V_{2}}E(V)=0.$ (81)
-Note that Eq.
-(80) and the boundary conditions (81) are symmetric with respect to interchange of the indexes $2$ and $1$, while Eq.
-(63) and the boundary conditions (49a) - certainly not.
-So to the boundary conditions (64) we should add the demand (61) (or opposite inequality for $V=V_{1}$).
-Substituting (61) into (80) we obtain the demand written in the new variables as
-$\left.\frac{T^{2}}{2}\frac{d^{2}E}{dV^{2}}\right|_{V=V_{2}}+\left.\tau_{R}\frac{dE}{dV}\right|_{V=V_{2}}<0.$ (82)
-Since (80) is the first order differential equation, the problem is overdetermined.
-However, for the problems considered explicitly in Section V.2.1, not only will we be able to find the solutions of (80) satisfying the boundary conditions (81) - the solutions will turn out to be simple polynomial (fractional polynomial) functions.
-For $Q(V)$ given by (71), Eq. (80) can be presented as
-$T^{2}E\frac{dE}{dV}+\tau_{R}E=\gamma(V-V_{1})(V-V_{2})(V+V_{3}).$ (83)
-The boundary conditions (81) and the r.h.s.
-of (83) prompt the idea to look for $E$ as the polynomial function proportional to $\left(V-V_{1}\right)\left(V-V_{2}\right)$.
-Because the r.h.s.
-of (83) is the third order polynomial of $V$, the multiplying factor should be a constant.
-That is the natural inspired guess is kogan4 ; kogan5
-$E(V)=\frac{\psi}{V_{1}-V_{2}}\left(V-V_{1}\right)\left(V-V_{2}\right)$ (84)
-(the normalization was chosen for later convenience).
-Substituting (84) into (83) we obtain
-$\frac{T^{2}\psi^{2}}{\left(V_{1}-V_{2}\right)^{2}}\left(2V-V_{1}-V_{2}\right)+\frac{\tau_{R}\psi}{V_{1}-V_{2}}=\gamma\left(V+V_{3}\right).$
-Hence the condition on the parameters of the problem, for (84) to be the solution of (83) is given by (73), and the constant multiplier in (84) should is given by (75).
-Substituting (84) into the equation
-$\frac{dV}{d\tau}=E(V)$ (86)
-we obtain the solution (74).
-For $Q(V)$ given by (66) we obtain, instead of (83), Abel equation of the second kind polyanin
-$T^{2}E\frac{dE}{dV}+\tau_{R}E=\beta(V-V_{1})(V-V_{2}).$ (87)
-Counting the powers of $V$ in the r.h.s.
-of (87) we come to the conclusion that if $E$ is a polynomial function of $V$, it should be a 3/2 order polynomial function.
-And the boundary conditions (81) tell us that this polynomial should zero out at $V=V_{1,2}$.
-To acquire an inspired guess how the polynomial may look like, let us return to the solution of (83) we obtained previously and split in two one multiplier in the solution
-$\frac{V-V_{1}}{V_{1}-V_{2}}=\left[\left(\frac{V-V_{2}}{V_{1}-V_{2}}\right)^{1/2}-1\right]\left[\left(\frac{V-V_{2}}{V_{1}-V_{2}}\right)^{1/2}+1\right].$
-All that was said above being taken together, prompts us to check up the solution of (87) in the form kogan4 ; kogan5
-$E(V)=2\chi\left[\left(\frac{V-V_{2}}{V_{1}-V_{2}}\right)^{1/2}-1\right]\left(V-V_{2}\right).$ (89)
-Substituting the ansatz into (87) we see that it is in fact the solution of the equation, provided the parameters of the problem satisfy the relation (68) and $\chi$ is given by (70).
-Substituting (89) into (86) we obtain (69)
-Actually, because of the symmetry of Eq.
-(80) and the boundary conditions (81) with respect to interchange of the indexes $2$ and $1$, instead of (89) and the condition (68) we could as well have obtained those with the indexes $2$ and $1$ interchanged.
-However, one can easily see that (89) satisfies the demand (82), and (89) with the indexes 1 and 2 interchanged - doesn’t.
+$R_{L}R_{C}\frac{dQ}{d\Phi}=\frac{U(\Phi-\Phi_{1})-V(Q)+V(Q_{1})}{U(Q-Q_{1})-I(\Phi)+I(\Phi_{1})}.$ (96)
+If we manage to solve (96), we can substitute the solution into one of the equations (93) and integrate the resulting equation in quadratures to obtain the profile of the shock wave.
+It is reassuring to realize that the conditions for the integrability in elementary functions of Eqs.
+(64) and (70) describing weak shocks in half-nonlinear transmission line (i.e.
+Eqs.
+(66) and (71)) are actually the (nonlinear) conditions on the product $R_{L}R_{C}$.
+Further analysis of (96) we postpone until later.

### v4

@@
 Abstract
-We consider interaction between the small amplitude travelling waves (”sound”) and the shock waves in the transmission line containing both nonlinear capacitors and nonlinear inductors.
-We calculate for the ”sound” wave the coefficient of reflection from (the coefficient of transmission through) the shock wave.
-These coefficients are expressed in terms of the wave speeds and the wave impedances.
-When only the capacitors or only the inductors are nonlinear, the coefficients are expressed in terms of the wave speeds only.
-We explicitly include into consideration of the shocks the dissipation, introducing ohmic resistors shunting the inductors and also in series with the capacitors.
-This allows us to describe the shocks as physical objects of finite width and study their profiles.
+In the first half of the paper we consider interaction between the small amplitude travelling waves (”sound”) and the shock waves in the transmission line containing both nonlinear capacitors and nonlinear inductors.
+We calculate the ”sound” wave coefficient of reflection from (coefficient of transmission through) the shock wave.
+These coefficients are expressed in terms of the speeds of the ”sound” waves relative to the shock and the wave impedances.
+In the second half of the paper we explicitly include into consideration the dissipation in the system, introducing ohmic resistors shunting the inductors and also in series with the capacitors.
+This allows us to justify the conditions on the shocks, postulated in the first half of the paper.
+This also allows us to describe the shocks as physical objects of finite width and study their profiles.
 In some particular cases the profiles were obtained in terms of elementary functions.
@@
 In the second part of the paper we include explicitly the ohmic losses (which are of course necessary for the existence of the shock waves) into the circuit equations.
-This allows us to study the profile of the shock wave.
 II The circuit equations
@@
 Figure 1: Lossless nonlinear transmission line
-We take the voltages on the capacitors $V_{n}$ and the current through the inductors $I_{n}$ as the dynamical variables. The circuit equations (Kirchhoff laws) are
+We take the capacitors charges $Q_{n}$ and the integrated voltages on the inductors
+[…]
+as the dynamical variables. The circuit equations (Kirchhoff laws) are
 $\frac{dQ_{n}}{dt}$ $=I_{n}-I_{n+1},$ (2a)
-$\frac{d\Phi_{n+1}}{dt}$ $=V_{n}-V_{n+1},$ (1b)
-where
-[…]
-$Q_{n}$ is the charge of the nonlinear capacitor and ${\cal V}_{n}$ is the voltage on the nonlinear inductor.
-In the continuum approximation we treat $n$ as the continuous variable $x$ and approximate the finite differences in the r.h.s.
+$\frac{d\Phi_{n+1}}{dt}$ $=V_{n}-V_{n+1}.$ (2b)
+To close the system (2) we should specify the connection between the voltages on the capacitors $V_{n}$ and the charges and also between the currents through the inductors $I_{n}$ and the integrated voltages
+$V_{n}$ $=V(Q_{n}),$ (3a)
+$I_{n}$ $=I(\Phi_{n}).$ (3b)
+Further on we’ll consider $V(Q)$ and $I(\Phi)$ as known functions.
+In the continuum approximation we treat $n$ as the continuous variable $x$ (we measure distance in the units of the transmission line period) and approximate the finite differences in the r.h.s.
 of the equations by the first derivatives with respect to $x$, after which the equations take the form
 $\frac{\partial Q}{\partial t}$ $=-\frac{\partial I}{\partial x},$ (4a)
-$\frac{\partial\Phi}{\partial t}$ $=-\frac{\partial V}{\partial x}$ (3b)
-(we measure distance in the units of the transmission line period). To close the system (3) we should specify
-$V$ $=V(Q),$ (4a)
-$I$ $=\Phi(\Phi).$ (4b)
-Further on we’ll consider $V$ and $I$ as known functions of $Q$ and $\Phi$ respectively.
+$\frac{\partial\Phi}{\partial t}$ $=-\frac{\partial V}{\partial x}.$ (4b)
 III The ”sound” and the shocks
@@
 In this paper we’ll be interested in only two types of those.
-First type - small amplitude waves on a homogeneous background $\varphi_{0},V_{0}$.
-$V$ $=V_{0}+v,$ (5a)
-$I$ $=I_{0}+i.$ (5b)
+First type - small amplitude waves on a homogeneous background $Q_{0},\Phi_{0}$.
+$Q$ $=Q_{0}+q,$ (5a)
+$\Phi$ $=\Phi_{0}+\phi.$ (5b)
 For such waves Eq. (4) is simplified to
-$C_{0}\frac{\partial v}{\partial t}$ $=-\frac{\partial i}{\partial x},$ (6a)
-$L_{0}\frac{\partial i}{\partial t}$ $=-\frac{\partial v}{\partial x},$ (6b)
+$\frac{\partial q}{\partial t}$ $=-\frac{1}{L_{0}}\frac{\partial\phi}{\partial x},$ (6a)
+$\frac{\partial\phi}{\partial t}$ $=-\frac{1}{C_{0}}\frac{\partial q}{\partial x},$ (6b)
 where
-$C_{0}$ $=\left.\frac{dQ}{dV}\right|_{Q=Q_{0}},$ (7a)
-$L_{0}$ $=\left.\frac{d\Phi}{dI}\right|_{\Phi=\Phi_{0}}.$ (7b)
-The solutions of Eq. (6) (which, for brevity we’ll call the ”sound”) are right- and left-propagating travelling waves, each depending upon the single variable $\tau_{\pm}=t\mp x/u_{0}$, where
-$u_{0}^{2}=\frac{1}{L_{0}C_{0}}$ (8)
-is the ”sound” speed. The current and the voltage in the waves are connected by the equations
-$v=\pm Z_{0}i,$ (9)
+[…]
+The solutions of Eq.
+(6) (which, for brevity we’ll call the ”sound”) are right- and left-propagating travelling waves, each depending upon the single variable $\tau_{\pm}=t\mp x/u_{0}$.
+The variables in the ”sound” wave are connected by the equations
+$v$ =±$=\pm u_{0}\phi,$ (8a)
+$i$ =±$=\pm u_{0}q,$ (8b)
+$\phi$ =±$=\pm Z_{0}q,$ (8c)
 where
-$Z_{0}=\sqrt{\frac{L_{0}}{C_{0}}}$ (10)
-is the linear impedance of the transmission line.
+$u_{0}$ $=\frac{1}{\sqrt{L_{0}C_{0}}},$ (9a)
+$Z_{0}$ $=\sqrt{\frac{L_{0}}{C_{0}}}$ (9b)
+are the ”sound” speed and the linear impedance of the transmission line respectively.
 The second type of solutions, we are interested in, is shock waves, containing (in the approximation of the present Section) discontinuities, satisfying the conditions (see Section V)
@@
 $U_{21}\left(\Phi_{2}-\Phi_{1}\right)$ $=V_{2}-V_{1},$ (10b)
-where
-$V_{a,b}\equiv V(Q_{a,b}),\hskip 28.45274ptI_{a,b}\equiv I(\Phi_{a,b}),$ (12)
-and $U_{21}$ is the shock wave speed.
+where $V_{a,b}\equiv V(Q_{a,b})$, $I_{a,b}\equiv I(\Phi_{a,b})$, and $U_{21}$ is the shock wave speed.
 The indices 1 and 2, which refer to the quantities before and after the shock respectively, enter in the system (10) in a symmetrical way.
@@
 For the first problem mentioned above we have
-$\left(\begin{array}[]{c|c}V_{1}&V_{2}\\ I_{1}&I_{2}\end{array}\right)=\left(\begin{array}[]{c|c}V_{b}&V_{a}+v^{(in)}+v^{(r)}\\ I_{b}&I_{a}+i^{(in)}+i^{(r)}\end{array}\right),$ (18)
+$\left(\begin{array}[]{c|c}Q_{1}&Q_{2}\\ \Phi_{1}&\Phi_{2}\end{array}\right)=\left(\begin{array}[]{c|c}Q_{b}&Q_{a}+q^{(in)}+q^{(r)}\\ \Phi_{b}&\Phi_{a}+\phi^{(in)}+\phi^{(r)}\end{array}\right),$ (16)
 where (in) stands for the incident and (r) - for the reflected ”sound” wave. We also have to take into account the change of the shock speed due to the ”sound” waves
-$U_{21}=U_{ab}+\delta U^{(refl)}.$ (19)
-Substituting (18) into (11), in linear with respect to the ”sound” wave amplitudes approximation we obtain
-$\delta U^{(refl)}\left(Q_{a}-Q_{b}\right)+U_{ab}C_{a}\left(v^{(in)}+v^{(r)}\right)=i^{(in)}+i^{(r)},$ (20a)
-$\delta U^{(refl)}\left(\Phi_{a}-\Phi_{b}\right)+U_{ab}L_{a}\left(i^{(in)}+i^{(r)}\right)=v^{(in)}+v^{(r)}.$ (20b)
-Dividing the equations (20) to each other we obtain
-$u_{in}v^{(in)}-u_{r}v^{(r)}=Z_{ab}\left(u_{in}i^{(in)}-u_{r}i^{(r)}\right),$ (21)
+$U_{21}=U_{ab}+\delta U$ (17)
+(same in the next problem). Substituting (16) into (10), in linear with respect to the ”sound” wave amplitudes approximation we obtain (taking into account (8a) and (8b))
+$\delta U\left(Q_{a}-Q_{b}\right)+U_{ab}\left(q^{(r)}+q^{(in)}\right)=u_{a}\left(q^{(in)}-q^{(r)}\right),$ (18a)
+$\delta U\left(\Phi_{a}-\Phi_{b}\right)+U_{ab}\left(\phi^{(r)}+\phi^{(in)}\right)=u_{a}\left(\phi^{(in)}-\phi^{(r)}\right).$ (18b)
+Excluding $\delta U$ from (18) and taking into account (8c) we obtain for the reflection coefficient
+$R\equiv\frac{\phi^{(r)}}{\phi^{(in)}}=\frac{\left(Z_{a}-Z_{ab}\right)u_{in}}{\left(Z_{a}+Z_{ab}\right)u_{r}},$ (19)
 where $u_{in}=u_{a}-U_{ab}$ is the speed of the incident ”sound” wave relative to the shock wave, $u_{r}=u_{a}+U_{ab}$ is the speed of the reflected ”sound” wave relative to the shock wave and we introduced the impedance of the shock wave
-$Z_{ab}\equiv\frac{V_{a}-V_{b}}{I_{a}-I_{b}}=\frac{\Phi_{a}-\Phi_{b}}{Q_{a}-Q_{b}}.$ (22)
-Taking into account the relations
-$v^{(in)}$ $=Z_{a}i^{(in)},$ (23a)
-$v^{(r)}$ $=-Z_{a}i^{(r)}$ (23b)
-(the difference in the signs is because of the opposite directions of propagation of the two waves) and solving (21) we obtain for the reflection coefficient
-$R\equiv\frac{i^{(r)}}{i^{(in)}}=\frac{\left(Z_{ab}-Z_{a}\right)u_{in}}{\left(Z_{ab}+Z_{a}\right)u_{r}}.$ (24)
+$Z_{ab}\equiv\frac{\Phi_{a}-\Phi_{b}}{Q_{a}-Q_{b}}.$ (20)
+Note that from (10) follows
+$\frac{\Phi_{a}-\Phi_{b}}{Q_{a}-Q_{b}}=\frac{V_{a}-V_{b}}{I_{a}-I_{b}}.$ (21)
 For the second problem we have
-$\left(\begin{array}[]{c|c}V_{1}&V_{2}\\ I_{1}&I_{2}\end{array}\right)=\left(\begin{array}[]{c|c}V_{b}+V^{(in)}&V_{a}+v^{(t)}\\ I_{b}+i^{(in)}&I_{a}+i^{(t)}\end{array}\right),$ (29)
+$\left(\begin{array}[]{c|c}Q_{1}&Q_{2}\\ \Phi_{1}&\Phi_{2}\end{array}\right)=\left(\begin{array}[]{c|c}Q_{b}+q^{(in)}&Q_{a}+q^{(t)}\\ \Phi_{b}+\phi^{(in)}&\Phi_{a}+\phi^{(t)}\end{array}\right),$ (26)
 where (t) stands for the transmitted wave. Substituting (26) into (10), in linear approximation we obtain
-$\delta U^{(trans)}\left(Q_{a}-Q_{b}\right)+U_{ab}\left(C_{a}v^{(t)}-C_{b}v^{(in)}\right)=i^{(t)}-i^{(in)},$ (30a)
-$\delta U^{(trans)}\left(\Phi_{a}-\Phi_{b}\right)+U_{ab}\left(L_{a}i^{(t)}-L_{b}i^{(in)}\right)=v^{(t)}-v^{(in)}.$ (30b)
-Dividing the equations (30) to each other we obtain
-$u_{a}u_{in}v^{(in)}-u_{b}u_{t}v^{(t)}=Z_{ab}\left(u_{a}u_{in}i^{(in)}-u_{b}u_{t}i^{(t)}\right),$
+$\delta U\left(Q_{a}-Q_{b}\right)+U_{ab}\left(q^{(t)}-q^{(in)}\right)=u_{b}q^{(in)}-u_{a}q^{(t)},$ (27a)
+$\delta U\left(\Phi_{a}-\Phi_{b}\right)+U_{ab}\left(\phi^{(t)}-\phi^{(in)}\right)=u_{b}\phi^{(in)}-u_{a}\phi^{(t)}.$ (27b)
+Excluding $\delta U$ from (27) and taking into account (8c) we obtain for the transmission coefficient
+$T\equiv\frac{\phi^{(t)}}{\phi^{(in)}}=\frac{\left(Z_{b}+Z_{ab}\right)Z_{a}u_{in}}{\left(Z_{a}+Z_{ab}\right)Z_{b}u_{t}},$ (28)
 where $u_{in}=u_{b}+U_{ab}$ is the speed of the incident ”sound” wave relative to the shock wave, and $u_{t}=u_{a}+U_{ab}$ is the speed of the transmitted ”sound” wave relative to the shock wave.
-Taking into account the relations
-$v^{(in)}$ $=-Z_{b}i^{(in)},$ (32a)
-$v^{(t)}$ $=-Z_{a}i^{(t)}$ (32b)
-and solving (IV) we obtain for the transmission coefficient
-$T\equiv\frac{i^{(t)}}{i^{(in)}}=\frac{\left(Z_{ab}+Z_{b}\right)u_{a}u_{in}}{\left(Z_{ab}+Z_{a}\right)u_{b}u_{t}}.$ (33)
 As one could have expected, the modulus of the ”sound” reflection coefficient is less than one, and it goes to zero when the intensity of the shock wave decreases, that is when $Q_{a}\to Q_{b},\Phi_{a}\to\Phi_{b}$, in other words, when the shock wave itself is nearly the ”sound” wave.
 Similarly, in these circumstances the transmission coefficient goes to 1.
-Looking back at (21) and (IV) we wonder whether these equations could have been written down by considering propagation of the waves in the frame of reference moving with the shock wave.
+Looking at (19) and (28) we wonder whether these equations could have been obtained by considering propagation of the waves in the frame of reference moving with the shock wave.
 IV.1 Half-nonlinear transmission line
 IV.1.1 Linear capacitor
-For linear capacitor ($Q=VC$, where $C$ is a constant) from (8) and (10) follows that
-$Z_{a,b}=\frac{1}{Cu_{a,b}},$ (34)
-and from (11) and (22) follows that
-$Z_{ab}=\frac{1}{CU_{ab}}.$ (35)
+For linear capacitor ($V=Q/C$, where $C$ is a constant)
+$Z_{a,b}=\frac{1}{Cu_{a,b}},\hskip 28.45274ptZ_{ab}=\frac{1}{CU_{ab}}.$ (29)
 Hence Eqs. (19) and (28) can be simplified respectively to
-$R$ $=\frac{u_{in}^{2}}{u_{r}^{2}},$ (36a)
-$T$ $=\frac{u_{a}^{2}u_{in}^{2}}{u_{b}^{2}u_{t}^{2}},$ (36b)
-and from (8) and (11) follows
+$R$ $=-\frac{u_{in}^{2}}{u_{r}^{2}},$ (30a)
+$T$ $=\frac{u_{in}^{2}}{u_{t}^{2}},$ (30b)
+and from (9a) and (10) follows
 $u_{a,b}^{2}$ $=\frac{1}{C}\left.\frac{dI}{d\Phi}\right|_{\Phi=\Phi_{a,b}}$ (31a)
 $U_{ab}^{2}$ $=\frac{1}{C}\frac{I_{a}-I_{b}}{\Phi_{a}-\Phi_{b}}.$ (31b)
-Consider the case of JTL kogan4 . In this case from the second Josephson law
-${\cal V}=\frac{\hbar}{2e}\frac{d\varphi}{dt},$ (38)
-where $\varphi$ is Josephson phase, we obtain
-$\Phi=\frac{\hbar}{2e}\varphi.$ (39)
-Hence the first Josephson law gives
+For example, for the JTL
 $I=I_{c}\sin\left(\frac{2e\Phi}{\hbar}\right),$ (32)
-and we can easily calculate all the velocities.
-Equations (36) for the case of JTL were obtained in our previous publication kogan4 .
-(Comparing the results of Ref.
-kogan4 with those presented in this paper, one should pay attention to a slightly different definitions of the transmission and the reflection coefficients in the two papers.)
+and all the velocities can be easily calculated. Equations (30) in this case were obtained in our previous publication kogan4 .
 IV.1.2 Linear inductor
-For linear inductor ($\Phi=LI$, where $L=$ is a constant) from (8) and (10) follows that
-$Z_{a,b}=Lu_{a,b},$ (41)
-and from (11) and (22) follows that
-$Z_{ab}=LU_{ab}.$ (42)
+For linear inductor ($I=\Phi/L$, where $L$ is a constant)
+$Z_{a,b}=Lu_{a,b},\hskip 28.45274ptZ_{ab}=LU_{ab}.$ (33)
 Hence Eqs. (19) and (28) can be respectively simplified to
-$R$ $=-\frac{u_{in}^{2}}{u_{r}^{2}},$ (43a)
-$T$ $=\frac{u_{a}u_{in}^{2}}{u_{b}u_{t}^{2}},$ (43b)
-and from (8) and (11) follows
+$R$ $=\frac{u_{in}^{2}}{u_{r}^{2}},$ (34a)
+$T$ $=\frac{Z_{a}u_{in}^{2}}{Z_{b}u_{t}^{2}},$ (34b)
+and from (9a) and (10) follows
 $u_{a,b}^{2}$ $=\frac{1}{L}\left.\frac{dV}{dQ}\right|_{Q=Q_{a,b}}$ (35a)
 $U_{ab}^{2}$ $=\frac{1}{L}\frac{V_{a}-V_{b}}{Q_{a}-Q_{b}}.$ (35b)
-Note that because of the symmetry of Eq. (3) with respect to simultaneous interchange
-$Q$ $\leftrightarrows\Phi,$ (45a)
-$V$ $\leftrightarrows I,$ (45b)
-Eqs. (43), (44) follow from Eqs. (36), (37).
 V Lossy transmission line
@@
 $\lim_{\tau\to-\infty}\Phi(\tau)$ $=\Phi_{1};\hskip 14.22636pt\lim_{\tau\to+\infty}\Phi(\tau)=\Phi_{2}.$ (39b)
-From the five parameters ($Q_{1},Q_{2},\Phi_{1},\Phi_{2},U$) - three are arbitrary constants, and the other two are defined by the equations.
-Both $(Q_{1},\Phi_{1})$ and $(Q_{2},\Phi_{2})$ are the fixed points (in the $(Q,\Phi)$ phase space), hence the r.h.s.
-of (48b) is equal to zero in both points.
-Equating these two values we recover the jump conditions (11b).
-Doing the same procedure with (48a) we recover the jump conditions (11a).
-$(Q_{1},\Phi_{1})$ is the semistable fixed point of the system (48), and $(Q_{2},\Phi_{2})$ is the stable fixed point. (Other options would have demanded fine tuning of the parameters.) This means
+Equating the r.h.s.
+of (38a) at $\tau=\pm\infty$, we recover the jump conditions (10a).
+Doing the same procedure with (38b) we recover the jump conditions (10b).
+Because the quantities ($Q_{1},Q_{2},\Phi_{1},\Phi_{2},U$), satisfy the constraint (21), Eq.
+(39) specifies 3 boundary conditions, which thus define also the shock speed $U$.
+Continuing the analysis of the boundary conditions, we notice that $(Q_{1},\Phi_{1})$ is the semistable fixed point of the system (38), and $(Q_{2},\Phi_{2})$ is the stable fixed point.
+(Other options would have demanded fine tuning of the parameters.) This means
 $\left|\begin{array}[]{cc}\frac{d}{d\Phi}\left[U_{21}Q-I(\Phi)\right]_{Q_{1},\Phi_{1}}&\frac{d}{dQ}\left[U_{21}Q-I(\Phi)\right]_{Q_{1},\Phi_{1}}\\ \frac{d}{d\Phi}\left[U_{21}\Phi-V(Q)\right]_{Q_{1},\Phi_{1}}&\frac{d}{dQ}\left[U_{21}\Phi-V(Q)\right]_{Q_{1},\Phi_{1}}\end{array}\right|<0,$ (42)
@@
 Because of the symmetry of Eq.
-(48) given by (45) (to which we should add $1/R_{L}\leftrightarrows R_{C}$), the case of the transmission line with linear capacitors and nonlinear inductors (JTL first and foremost) would be identical to the one considered in this Section..
-Because of the relation $\Phi=LI$, where $L$ is a constant, we can exclude the current from (48) and after a bit of algebra obtain
+(38) the case of the transmission line with linear capacitors and nonlinear inductors (JTL first and foremost) would be identical to the one considered in this Section..
+We can exclude the current from (38) and after a bit of algebra obtain
 $\frac{LR_{C}}{R_{L}}\frac{d^{2}Q}{d\tau^{2}}+R_{C}\frac{dQ}{d\tau}+\frac{L}{R_{L}}\frac{dV(Q)}{d\tau}=F(Q),$ (49)
@@
 We are also grateful to the organizers of Nor-Amberd School in Theoretical Physics, the scientific atmosphere of which stimulated the presented research.
-The paper was finalized during the Author’s participation in School and Workshop on Dynamical Systems, ICTP, Trieste.
+The paper was finalized during the Author’s participation in School and Workshop on Dynamical Systems and Workshop on Localization and Ergodicity, ICTP, Trieste.
 The Author is grateful to the Center for the hospitality.
@@
 Substituting (68) into Eq.
-(74) we obtain the solution (72).
+(64) and integrating we obtain the solution (62).
 For $Q(V)$ given by (53) we obtain, instead of (67),
 $T^{2}{\cal I}\frac{d{\cal I}}{dQ}+\tau_{R}{\cal I}=\beta(Q-Q_{1})(Q-Q_{2}).$ (70)
-If ${\cal I}$ is a polynomial function of $Q$, it should be a 3/2 order polynomial. And the boundary conditions (76) tell us that this polynomial should zero out at $Q=Q_{1,2}$.
-Let us split in two one multiplier in the solution (78)
+If ${\cal I}$ is a polynomial function of $Q$, it should be a 3/2 order polynomial equal to zero at $Q=Q_{1,2}$. So let us split in two one multiplier in the solution (68)
 $\frac{Q-Q_{1}}{Q_{1}-Q_{2}}=\left[\left(\frac{Q-Q_{2}}{Q_{1}-Q_{2}}\right)^{1/2}-1\right]\left[\left(\frac{Q-Q_{2}}{Q_{1}-Q_{2}}\right)^{1/2}+1\right].$
@@
 Substituting the ansatz into (70) we see that it is in fact the solution of the equation, provided the parameters of the problem satisfy the relation (56) and $\chi$ is given by (58).
-Substituting thus obtained ${\cal I}(Q)$ into (74) we obtain (67)
+Substituting thus obtained ${\cal I}(Q)$ into (64) and integrating we obtain (57)
 A.2 Solitary wave solution of a generalized KdV equation

### v5

@@
 This allows us to justify the conditions on the shocks, postulated in the first half of the paper.
-This also allows us to describe the shocks as physical objects of finite width and study their profiles.
-In some particular cases the profiles were obtained in terms of elementary functions.
+This also allows us to describe the shocks as physical objects of finite width and study their profiles, same as the profiles of the waves closely connected with the shocks - the kinks.
+The profiles of the latter, and in some particular cases the profiles of the former, were obtained in terms of elementary functions.
 I Introduction
@@
 For such waves (37) turns into the system of ODE which after integration takes the form
-$\frac{1}{R_{L}}\frac{d\Phi}{d\tau}$ $=UQ-I(\Phi)+\text{const},$ (38a)
-$R_{C}\frac{dQ}{d\tau}$ $=U\Phi-V(Q)+\text{const},$ (38b)
-To study the shock waves we should impose the boundary conditions
+$\frac{1}{R_{L}}\frac{d\Phi}{d\tau}$ $=UQ-I(\Phi)+\text{const}_{1},$ (38a)
+$R_{C}\frac{dQ}{d\tau}$ $=U\Phi-V(Q)+\text{const}_{2},$ (38b)
+Equation (38) contains 3 arbitrary constants - $U,\text{const}_{1},\text{const}_{2}$ hence to define the solution unequivocally we need 5 boundary conditions.
+Because we are ready to consider the solution up to translation in time, 4 of those would be enough.
+Since we are considering localized travelling waves, it would be convenient to take as those the quantities $Q_{1},Q_{2},\Phi_{1},\Phi_{2}$, defined by the equations
 $\lim_{\tau\to-\infty}Q(\tau)$ $=Q_{1};\hskip 14.22636pt\lim_{\tau\to+\infty}Q(\tau)=Q_{2};$ (39a)
 $\lim_{\tau\to-\infty}\Phi(\tau)$ $=\Phi_{1};\hskip 14.22636pt\lim_{\tau\to+\infty}\Phi(\tau)=\Phi_{2}.$ (39b)
-Equating the r.h.s.
-of (38a) at $\tau=\pm\infty$, we recover the jump conditions (10a).
-Doing the same procedure with (38b) we recover the jump conditions (10b).
-Because the quantities ($Q_{1},Q_{2},\Phi_{1},\Phi_{2},U$), satisfy the constraint (21), Eq.
-(39) specifies 3 boundary conditions, which thus define also the shock speed $U$.
-Continuing the analysis of the boundary conditions, we notice that $(Q_{1},\Phi_{1})$ is the semistable fixed point of the system (38), and $(Q_{2},\Phi_{2})$ is the stable fixed point.
+Equating the r.h.s. of (38a) at $\tau=\pm\infty$, we recover the jump conditions (10a). Doing the same with (38b) we recover the jump conditions (10b).
+The boundary conditions show that $(Q_{1},\Phi_{1})$ is the semistable fixed point of the system (38), and $(Q_{2},\Phi_{2})$ is the stable fixed point (see, however, Section VIII.1).
 (Other options would have demanded fine tuning of the parameters.) This means
-$\left|\begin{array}[]{cc}\frac{d}{d\Phi}\left[U_{21}Q-I(\Phi)\right]_{Q_{1},\Phi_{1}}&\frac{d}{dQ}\left[U_{21}Q-I(\Phi)\right]_{Q_{1},\Phi_{1}}\\ \frac{d}{d\Phi}\left[U_{21}\Phi-V(Q)\right]_{Q_{1},\Phi_{1}}&\frac{d}{dQ}\left[U_{21}\Phi-V(Q)\right]_{Q_{1},\Phi_{1}}\end{array}\right|<0,$ (42)
+$\left|\begin{array}[]{cc}\frac{d}{d\Phi}\left[U_{21}Q-I(\Phi)\right]_{Q_{2},\Phi_{2}}&\frac{d}{dQ}\left[U_{21}Q-I(\Phi)\right]_{Q_{2},\Phi_{2}}\\ \frac{d}{d\Phi}\left[U_{21}\Phi-V(Q)\right]_{Q_{2},\Phi_{2}}&\frac{d}{dQ}\left[U_{21}\Phi-V(Q)\right]_{Q_{2},\Phi_{2}}\end{array}\right|>0,$ (42)
 or
-$\left|\begin{array}[]{cc}-\left.\frac{dI}{d\Phi}\right|_{\Phi_{1}}&U_{21}\\ U_{21}&-\left.\frac{dV}{dQ}\right|_{Q_{1}}\end{array}\right|<0.$ (46)
+$\left|\begin{array}[]{cc}-\left.\frac{dI}{d\Phi}\right|_{\Phi_{2}}&U_{21}\\ U_{21}&-\left.\frac{dV}{dQ}\right|_{Q_{2}}\end{array}\right|>0.$ (46)
 Calculating the determinant and recalling the result for the ”sound” speed we obtain
-$U_{21}^{2}>u_{1}^{2}.$ (47)
-Similarly for the point $(Q_{2},\Phi_{2})$ we obtain
-$u_{2}^{2}>U_{21}^{2}.$ (48)
+$U_{21}^{2}<u_{2}^{2}.$ (47)
+Similarly for the point $(Q_{1},\Phi_{1})$ we obtain
+$u_{1}^{2}<U_{21}^{2}.$ (48)
 Thus we proved the inequalities (11) postulated previously.
-Note that the kinks, which we considered in our previous publications in addition to shocks but do not consider here, correspond to both fixed points being semistable.
-Thus the kinks turn out to be supersonic both with respect to the phase before the front and the phase behind the front.
-V.1 Half-nonlinear transmission line
-Further analysis of the fully nonlinear line we relegate to Appendix B) and here consider transmission line with the linear capacitor.
+VI Half-nonlinear transmission line
+Further analysis of the fully nonlinear line we relegate to Appendix A) and here consider transmission line with the linear capacitor.
 Because of the symmetry of Eq.
-(38) the case of the transmission line with linear capacitors and nonlinear inductors (JTL first and foremost) would be identical to the one considered in this Section..
-We can exclude the current from (38) and after a bit of algebra obtain
+(38) the case of the transmission line with linear capacitors and nonlinear inductors (JTL first and foremost) would be identical to the one considered in this Section.
+We can exclude the current from (38) and after a bit of algebra obtain closed equation for $Q(\tau)$
 $\frac{LR_{C}}{R_{L}}\frac{d^{2}Q}{d\tau^{2}}+R_{C}\frac{dQ}{d\tau}+\frac{L}{R_{L}}\frac{dV(Q)}{d\tau}=F(Q),$ (49)
@@
 $F(Q)\equiv\frac{V_{2}-V_{1}}{Q_{2}-Q_{1}}Q-V(Q)+\frac{V_{1}Q_{2}-V_{2}Q_{1}}{Q_{2}-Q_{1}}.$ (50)
-To (49) we should add the boundary conditions (39a). (The third arbitrary constant appears when we connect $\Phi$ with the found $Q$.)
-Equation (49) describes motion with friction of a fictitious Newtonian particle, $Q$ being the coordinate of the particle, in the potential well defined by (50).
-The boundary conditions (39a) tell that the particle starts its motion at $\tau=-\infty$ at the potential maximum at $Q=Q_{1}$ and ends - at $\tau=+\infty$ at the potential minimum at $Q=Q_{2}$.
+The solution of (49) should satisfy the boundary conditions (39a). (The third constant appears when we connect $\Phi$ with the found $Q$.)
+Equation (49) describes motion with friction of a fictitious Newtonian particle, $Q$ being the coordinate of the particle, in the potential well $\Pi(Q)$ defined by the equation
+$-\frac{d\Pi(Q)}{dQ}=F(Q).$ (51)
+The boundary conditions (39) tell that the fictitious particle starts its motion at $\tau=-\infty$ at the potential maximum of $\Pi(Q)$ at $Q=Q_{1}$ and ends - at $\tau=+\infty$ at the potential minimum at $Q=Q_{2}$ (see, however, Section VIII.1).
 We may present these maximum/minimum conditions as
 $\left.\frac{dF(Q)}{dQ}\right|_{Q=Q_{2}}=\frac{V_{2}-V_{1}}{Q_{2}-Q_{1}}-\left.\frac{dV}{dQ}\right|_{Q=Q_{2}}<0$ (52)
-and opposite inequality for $V=V_{1}$. Recalling (35) we realise that we have reproduced in this particular case Eqs. (47), (48).
-V.1.1 Weak shocks
-Before going further it is convenient to rewrite (50) as
+and opposite inequality for $V=V_{1}$. Note that we have reproduced (46) in the particular case of half-nonlinear transmission line. It would be convenient further on to rewrite (50) as
 $F(Q)=\frac{\left[V(Q)-V_{1}\right]\left(Q-Q_{2}\right)-\left\{2\leftrightarrows 1\right\}}{Q_{2}-Q_{1}},$ (53)
 where the notation $\left\{2\leftrightarrows 1\right\}$ means the preceding term with the indexes $2$ and $1$ interchanged.
-For weak shocks, when $|Q_{1}-Q_{2}|$ is small, we can approximate $V(Q)$ for $Q$ between $Q_{1}$ and $Q_{2}$ as
+For weak shocks, when $|Q_{1}-Q_{2}|$ is small, we can approximated $V(Q)$ for $Q$ between $Q_{1}$ and $Q_{2}$ as
 $V(Q)=\text{const}+\frac{Q+\beta Q^{2}}{C_{s}},$ (54)
-where $\beta$ and $C_{s}$ are assumed to be constant. We also may take $V(Q)$ in the last term in the l.h.s. of (49) in the linear approximation. So finally Eq. (49) becomes
+where $\beta$ and $C_{s}$ are constants. We also took $V(Q)$ in the last term in the l.h.s. of (49) in the linear approximation. In this cases, (49) becomes
 $T^{2}\frac{d^{2}Q}{d\tau^{2}}+\tau_{R}\frac{dQ}{d\tau}=\beta(Q-Q_{1})(Q-Q_{2}),$ (55)
 where
-$\tau_{R}$ $\equiv R_{C}C_{s}+\frac{L}{R_{L}},$ (55a)
-$T^{2}$ $\equiv\frac{LC_{s}R_{C}}{R_{L}}.$ (55b)
-One can check up that when the parameters of the problem satisfy the condition
-$\frac{\tau_{R}^{2}}{T^{2}}=\frac{25\beta(Q_{1}-Q_{2})}{6},$ (56)
-there exists the elementary solution of (54)
+$T^{2}$ $\equiv\frac{LC_{s}R_{C}}{R_{L}},$ (56a)
+$\tau_{R}$ $\equiv R_{C}C_{s}+\frac{L}{R_{L}}.$ (56b)
+Note that
+$\Pi(Q_{1})-\Pi(Q_{2})=\frac{\beta}{6}\left(Q_{1}-Q_{2}\right)\left(Q_{1}^{2}-Q_{1}Q_{2}+Q_{2}^{2}\right),$
+hence
+$\beta(Q_{1}-Q_{2})>0.$ (58)
+We can also improve the approximation (54) by modifying it to
+$V(Q)=\text{const}+\frac{Q+\beta Q^{2}+\gamma Q^{3}}{C_{s}},$ (59)
+where $\gamma$ is also assumed to be constant.
+This modification can be important when $Q_{1},Q_{2}$ are close to zero and, as it normally happens, the charge is an odd function of the voltage.
+In this cases, (49) becomes
+$T^{2}\frac{d^{2}Q}{d\tau^{2}}+\tau_{R}\frac{dQ}{d\tau}=\gamma(Q-Q_{1})(Q-Q_{2})(Q+Q_{3}),$ (60)
+where
+$Q_{3}\equiv\frac{\beta}{\gamma}+Q_{1}+Q_{2}.$ (61)
+VII The ODE which doesn’t contain explicitly the independent variable
+We would be looking for the solutions of (55) and (60) in terms of elementary functions.
+Following the well known in mathematics principle, stating that the more general the problem is, the easier it is to solve it, let us change gears and instead of (60) consider the generalized damped Helmholtz-Duffing equation kogan4
+$x_{\tau\tau}+kx_{\tau}=\gamma x\left(x^{n}-x_{1}\right)\left(x^{n}+x_{3}\right),$ (62)
+where $n$, $k$, $\gamma$, $x_{1}$, $x_{3}$ are constants, with the boundary conditions
+$\lim_{\tau\to-\infty}x(\tau)=x_{1}^{1/n},\hskip 28.45274pt\lim_{\tau\to+\infty}x(\tau)=0.$ (63)
+Equation (62) doesn’t contain explicitly the independent variable $\tau$. This prompts the idea to consider $x$ as the new independent variable and
+$p=\frac{dx}{d\tau}$ (64)
+as the new dependent variable. In the new variables (62) takes the form of Abel equation of the second kind polyanin .
+$pp_{x}+kp=\gamma x\left(x^{n}-x_{1}\right)\left(x^{n}+x_{3}\right).$ (65)
+The boundary conditions in the new variables are
+$p\left(x_{1}^{1/n}\right)=p(0)=0.$ (66)
+One can easily check up that for $\gamma$ and $k$ connected by the formula
+$k^{2}=\frac{\gamma\left[x_{1}+(n+1)x_{3}\right]^{2}}{n+1},$ (67)
+the solution of (65) satisfying the boundary conditions (66) is
+$p=mx\left(x^{n}-x_{1}\right),\hskip 28.45274ptm=\sqrt{\frac{\gamma}{n+1}}.$ (68)
+Substituting $p(x)$ into (64) and integrating we obtain the solution of (62) as
+$x(\tau)=\frac{x_{1}^{1/n}.}{\left[\exp\left(nmx_{1}\tau\right)+1\right]^{1/n}}.$ (69)
+Now consider the equation
+$x_{\tau\tau}+kx_{\tau}=\beta x\left(x^{n}-x_{1}\right)$ (70)
+with the boundary conditions (63). We can present (70) as
+$x_{\tau\tau}+kx_{\tau}=\beta x\left(x^{n/2}-x_{1}^{1/2}\right)\left(x^{n/2}+x_{1}^{1/2}\right).$ (71)
+Hence we realize that for $k$ and $\beta$ connected by the formula
+$k=(n+4)\sqrt{\frac{\beta x_{1}}{2(n+2)}},$ (72)
+the solution of (70) satisfying the boundary conditions (66) is
+$x(\tau)=\frac{x_{1}^{1/n}.}{\left\{\exp\left[n\sqrt{\frac{\beta x_{1}}{2(n+2)}}\tau\right]+1\right\}^{2/n}}.$ (73)
+To additionally illustrate the method of integration used in this Section, in the Appendix B we recover in the framework of the method the solitary wave solution of a generalized KdV equation drazin
+Let us return to (62) and modify it to
+$x_{\tau\tau}+k(1+2\beta x^{n})x_{\tau}=\gamma x\left(x^{n}-x_{1}\right)\left(x^{n}+x_{3}\right).$ (74)
+Thus we take into account possible nonlinearity of friction. Thus instead of (65) we obtain
+$pp_{x}+k(1+2\beta x^{n})p=\gamma x\left(x^{n}-x_{1}\right)\left(x^{n}+x_{3}\right).$ (75)
+Acting as above we obtain that for $k$ and $\gamma$ connected by the formula
+$k^{2}=\frac{\gamma[x_{1}+(n+1)x_{3}]^{2}}{\left(1-2\beta x_{3}\right)\left(n+1+2\beta x_{1}\right)},$ (76)
+the solution of (75) satisfying the boundary conditions (66) is (68) (same as it was for $\beta=0$), only this time
+$m=\sqrt{\frac{\gamma\left(1-2\beta x_{3}\right)}{n+1+2\beta x_{1}}}.$ (77)
+Hence the solution of (74) with $\beta\neq 0$ is of the same form as for $\beta=0$ (Eq. (69)), only with the modified $m$.
+VIII Back to the transmission line
+Now let us return to the particular cases of Eq.
+(49) (with the boundary conditions (39a)) considered in Section VI.
+For (60) using the results of the previous Section we claim that when the boundary conditions satisfy the relation
+$\frac{\tau_{R}^{2}}{T^{2}}=\frac{\gamma\left(Q_{1}+Q_{2}+2Q_{3}\right)^{2}}{2},$ (78)
+the solution of the equation can be expressed in terms of elementary functions:
+$Q=Q_{2}+\frac{Q_{1}-Q_{2}}{\exp\left(\psi\tau\right)+1},$ (79)
+where
+$\psi=\sqrt{\frac{\gamma}{2}}\cdot\frac{Q_{1}-Q_{2}}{T}.$ (80)
+For (55) we claim that when the boundary conditions satisfy the relation
+$\frac{\tau_{R}}{T}=5\sqrt{\frac{\beta\left(Q_{1}-Q_{2}\right)}{6}},$ (81)
+the solution of the equation can be expressed in terms of elementary functions:
 $Q=Q_{2}+\frac{Q_{1}-Q_{2}}{\left[\exp\left(\chi\tau\right)+1\right]^{2}},$ (82)
 where
-$\chi=\frac{\tau_{R}}{5T^{2}}.$ (58)
-Approximation (53) can be inadequate, if $Q_{1},Q_{2}$ are close to zero and, as it normally happens, the charge is an odd function of the voltage. So we are motivated to generalize (53) to
-$V(Q)=\text{const}+\frac{Q+\beta Q^{2}+\gamma Q^{3}}{C_{s}}.$ (59)
-In this case (49) becomes
-$T^{2}\frac{d^{2}Q}{d\tau^{2}}+\tau_{R}\frac{dQ}{d\tau}=\gamma(Q-Q_{1})(Q-Q_{2})(Q+Q_{3}),$ (60)
-where $Q_{3}=\beta/\gamma+Q_{1}+Q_{2}$. One can check up that when the parameters of the problem satisfy the condition
-$\frac{\tau_{R}^{2}}{T^{2}}=\frac{\gamma\left(Q_{1}+Q_{2}+2Q_{3}\right)^{2}}{2},$ (61)
-there exists the elementary solution of (60)
-$Q=Q_{2}+\frac{Q_{1}-Q_{2}}{\exp\left(\psi\tau\right)+1},$ (62)
-where
-$\psi=\frac{\tau_{R}\left(Q_{1}-Q_{2}\right)}{T^{2}\left(Q_{1}+Q_{2}+2Q_{3}\right)}.$ (63)
-In the Appendix A we’ll show how one arrives to the solutions (57) and (62).
+$\chi=\sqrt{\frac{\beta(Q_{1}-Q_{2})}{6}}\cdot\frac{1}{T}.$ (83)
+Note that neither (78) goes to (81) nor (79) goes to (82) when $\gamma\to 0$.
+Small change in the equation leads to big change of the condition for the solution to be expressed in elementary functions.
+We used previously the linear approximation for $V(Q)$ in the last term in the l.h.s.
+of (49).
+Strictly speaking, since we are considering nonlinear $V(Q)$ in the r.h.s.
+of (49) more consistent would be to treat the same way the l.h.s.
+Equations.
+(76), (77) allow us to go one step in this direction, that is to consider $V(Q)$ in the last term in the l.h.s.
+of (49) in quadratic approximation (for cubic nonlinearity of $V(Q)$).
+The modification of Eqs.
+(78), (79) can be easily obtained from Eqs.
+(76), (77) for $n=1$.
+VIII.1 The shocks vs. the kinks
+The main subject of the present paper is the shock waves.
+However, Eq.
+(49) can describe both the shocks and the kinks kogan4 .
+To understand the difference between two types of waves note that, strictly speaking, the boundary conditions (39) tell us that the fictitious particle ends its motion at $\tau=+\infty$ at the stationary point $Q_{2}$, which can be either potential minimum or potential maximum.
+For former case corresponds to the shocks, the latter - to the kinks kogan4 .
+Equation (55) can describe only the shocks.
+Equation (60), when the point $-Q_{3}$ is between $Q_{2}$ and $Q_{1}$, describes the kink, when the point is outside of that interval - the shocks.
+Equations (78) and (79) are valid both for the kinks and for the shocks, but in these two cases the condition (78) has different meanings.
+For the kinks it’s the condition of their existence.
+For the shocks it’s the condition for them to be expressed in elementary functions (actually the condition that the fictitious particle doesn’t oscillate in the vicinity of the stable equilibrium).
+But, of course, the shocks exist for the general boundary conditions.
+The same can be said about the shocks described by Eq.
+(55).
 Acknowledgements.
@@
 Malomed, A.
-Novic-Cohen and V.
-Verbuk for the insightful discussions.
+Novic-Cohen, V.
+Verbuk and A.
+Volkov for the insightful discussions.
 We are also grateful to the organizers of Nor-Amberd School in Theoretical Physics, the scientific atmosphere of which stimulated the presented research.
@@
 The Author is grateful to the Center for the hospitality.
-Appendix A The ODE which doesn’t contain explicitly the independent variable
-Equation (49) doesn’t contain explicitly the independent variable $\tau$. This prompts the idea to consider $Q$ as the new independent variable and
-${\cal I}=\frac{dQ}{d\tau}$ (64)
-as the new dependent variable. In the new variables (49) takes the form of Abel equation of the second kind polyanin
-$T^{2}{\cal I}\frac{d{\cal I}}{dQ}+\tau_{R}{\cal I}=F(Q).$ (65)
-As it follows from (39a), the boundary conditions for (65) are
-$\lim_{Q\to Q_{1},Q_{2}}{\cal I}(Q)=0.$ (66)
-Since (65) is the first order differential equation, the problem is overdetermined.
-However, for the problems considered explicitly in Section V.1.1, not only will we be able to find the solutions of (65) satisfying the boundary conditions (66) - the solutions will turn out to be simple polynomial (fractional polynomial) functions.
-A.1 Integration in terms of elementary functions
-For $Q(V)$ given by (59), Eq. (65) can be presented as
-$T^{2}{\cal I}\frac{d{\cal I}}{dQ}+\tau_{R}{\cal I}=\gamma(Q-Q_{1})(Q-Q_{2})(Q+Q_{3}).$ (67)
-The boundary conditions (66) and the r.h.s.
-of (67) prompt the idea kogan4 ; kogan5 to look for ${\cal I}(Q)$ as the polynomial function proportional to $\left(Q-Q_{1}\right)\left(Q-Q_{2}\right)$.
-Because the r.h.s.
-of (67) is the third order polynomial of $V$, the multiplying factor should be a constant
-${\cal I}(Q)=\frac{\psi}{Q_{1}-Q_{2}}\left(Q-Q_{1}\right)\left(Q-Q_{2}\right)$ (68)
-(the normalization was chosen for later convenience).
-Substituting (68) into (67) we obtain
-$\frac{T^{2}\psi^{2}}{\left(Q_{1}-Q_{2}\right)^{2}}\left(2Q-Q_{1}-Q_{2}\right)+\frac{\tau_{R}\psi}{Q_{1}-Q_{2}}=\gamma\left(Q+Q_{3}\right).$
-Hence the condition on the parameters of the problem, for (68) to be the solution of (67) is given by (61), and the constant multiplier in (68) should is given by (63).
-Substituting (68) into Eq.
-(64) and integrating we obtain the solution (62).
-For $Q(V)$ given by (53) we obtain, instead of (67),
-$T^{2}{\cal I}\frac{d{\cal I}}{dQ}+\tau_{R}{\cal I}=\beta(Q-Q_{1})(Q-Q_{2}).$ (70)
-If ${\cal I}$ is a polynomial function of $Q$, it should be a 3/2 order polynomial equal to zero at $Q=Q_{1,2}$. So let us split in two one multiplier in the solution (68)
-$\frac{Q-Q_{1}}{Q_{1}-Q_{2}}=\left[\left(\frac{Q-Q_{2}}{Q_{1}-Q_{2}}\right)^{1/2}-1\right]\left[\left(\frac{Q-Q_{2}}{Q_{1}-Q_{2}}\right)^{1/2}+1\right].$
-This prompts us the idea to look up for the solution of (70) in the form kogan4 ; kogan5
-${\cal I}(Q)=2\chi\left[\left(\frac{Q-Q_{2}}{Q_{1}-Q_{2}}\right)^{1/2}-1\right]\left(Q-Q_{2}\right).$ (72)
-Substituting the ansatz into (70) we see that it is in fact the solution of the equation, provided the parameters of the problem satisfy the relation (56) and $\chi$ is given by (58).
-Substituting thus obtained ${\cal I}(Q)$ into (64) and integrating we obtain (57)
-A.2 Solitary wave solution of a generalized KdV equation
-To additionally illustrate the method of integration used in this Appendix, let us recover in the framework of the method the solitary wave solution of a generalized KdV equation drazin
+Appendix A Shock profile for the fully nonlinear transmission line
+Let us return to (38).
+The equation is simplified when there is a leading mechanism of dissipation, in which cases we should substitute into (38) either $R_{C}=0$ or $R_{L}=\infty$.
+In the first case we equate the r.h.s.
+of (38b) to zero and obtain
+$V=U\Phi+\text{const}_{2}.$ (84)
+Substituting $V$ into (3a), solving the equation with respect to $Q$ and then substituting thus obtained $Q(\Phi)$ into (38a) we obtain the equation which can be easily integrated in quadratures.
+In the second case we equate the r.h.s.
+of (38a) to zero and obtain
+$I=UQ+\text{const}_{1}.$ (85)
+Substituting $I$ into (3b), solving the equation with respect to $\Phi$ and then substituting thus obtained $\Phi(Q)$ into (38b) we obtain the equation which can be integrated in quadratures.
+Thus we obtained Eqs.
+(48a) and (48b) of Ref.
+kogan3 .
+If both $R_{C}>0$ and $R_{L}<\infty$, from (38) we may obtain the equation
+$R_{L}R_{C}\frac{dQ}{d\Phi}=\frac{U\Phi-V+\text{const}_{2}}{UQ-I+\text{const}_{1}}.$ (86)
+If we manage to solve (86), (3) we can substitute the solution into one of the equations (38) and integrate the resulting equation in quadratures to obtain the profile of the shock wave.
+It is reassuring to realize that the conditions for the integrability in elementary functions of Eqs.
+(55) and (60) describing weak shocks in half-nonlinear transmission line (i.e.
+Eqs.
+(78) and (81)) are actually the (nonlinear) conditions on the product $R_{L}R_{C}$.
+Further analysis of (86) we postpone until later.
+Appendix B Solitary wave solution of a generalized KdV equation
+Consider a generalized KdV equation drazin
 $\phi_{t}+(n+1)(n+2)\phi^{n}\phi_{x}+\phi_{xxx}=0.$ (87)
@@
 $\phi^{n}=\frac{U}{2}\operatorname{sech}^{2}\left[\frac{n\sqrt{U}}{2}\left(x-Ut\right)\right].$ (95)
-Appendix B Shock profile for the fully nonlinear transmission line
-For the fully nonlinear transmission line it is convenient to present the problem in the form slightly different from that used everywhere above.
-We’ll consider the shock speed $U$ and the conditions before the shock as given.
-Thus the boundary conditions are
-$\lim_{\tau\to-\infty}Q(\tau)=Q_{1};\hskip 14.22636pt\lim_{\tau\to-\infty}\Phi(\tau)$ $=\Phi_{1}$ (82)
-and we can concretize (38) to
-$\frac{1}{R_{L}}\frac{d\Phi}{d\tau}$ $=U(Q-Q_{1})-I+I(\Phi_{1}),$ (83a)
-$R_{C}\frac{dQ}{d\tau}$ $=U(\Phi-\Phi_{1})-V+V(Q_{1}).$ (83b)
-Equation (83) is simplified when there is a leading mechanism of dissipation, in which cases we should substitute into (83) either $R_{C}=0$ or $R_{L}=\infty$.
-In the first case we equate the r.h.s.
-of (83b) to zero and obtain
-$V=U(\Phi-\Phi_{1})+V(Q_{1}).$ (84)
-Substituting $V$ into (3a), solving the equation with respect to $Q$ and then substituting thus obtained $Q(\Phi)$ into (83a) we obtain the equation which can be easily integrated in quadratures.
-In the second case we equate the r.h.s.
-of (83a) to zero and obtain
-$I=U(Q-Q_{1})+(\Phi_{1}).$ (85)
-Substituting $I$ into (3b), solving the equation with respect to $\Phi$ and then substituting thus obtained $\Phi(Q)$ into (83b) we obtain the equation which can be integrated in quadratures.
-Thus we obtained Eqs.
-(48a) and (48b) of Ref.
-kogan3 .
-If both $R_{C}>0$ and $R_{L}<\infty$, from (83) we may obtain the equation
-$R_{L}R_{C}\frac{dQ}{d\Phi}=\frac{U(\Phi-\Phi_{1})-V(Q)+V(Q_{1})}{U(Q-Q_{1})-I(\Phi)+I(\Phi_{1})}.$ (86)
-If we manage to solve (86), we can substitute the solution into one of the equations (83) and integrate the resulting equation in quadratures to obtain the profile of the shock wave.
-It is reassuring to realize that the conditions for the integrability in elementary functions of Eqs.
-(54) and (60) describing weak shocks in half-nonlinear transmission line (i.e.
-Eqs.
-(56) and (61)) are actually the (nonlinear) conditions on the product $R_{L}R_{C}$.
-Further analysis of (86) we postpone until later.

### v6

@@
 where the notation $\left\{2\leftrightarrows 1\right\}$ means the preceding term with the indexes $2$ and $1$ interchanged.
-For weak shocks, when $|Q_{1}-Q_{2}|$ is small, we can approximated $V(Q)$ for $Q$ between $Q_{1}$ and $Q_{2}$ as
+Equation (49) can be integrated numerically, however in this paper we would be looking for analytical solutions.
+Trivially, for weak shocks, when $|Q_{1}-Q_{2}|$ is small, we can discard the term with the second derivative in the l.h.s.
+of the equation and it can be integrated in quadratures.
+However we’ll see that (49) can be integrated analytically (and even in terms of elementary functions) also in the case when both terms in the l.h.s.
+of the equation are of the same order of magnitude, though under additional assumptions.
+Let us approximate $V(Q)$ for $Q$ between $Q_{1}$ and $Q_{2}$ as
 $V(Q)=\text{const}+\frac{Q+\beta Q^{2}}{C_{s}},$ (54)
-where $\beta$ and $C_{s}$ are constants. We also took $V(Q)$ in the last term in the l.h.s. of (49) in the linear approximation. In this cases, (49) becomes
+where $\beta$ and $C_{s}$ are constants. Let us also take $V(Q)$ in the last term in the l.h.s. of (49) in the linear approximation. In this cases, (49) becomes
 $T^{2}\frac{d^{2}Q}{d\tau^{2}}+\tau_{R}\frac{dQ}{d\tau}=\beta(Q-Q_{1})(Q-Q_{2}),$ (55)
@@
 VII The ODE which doesn’t contain explicitly the independent variable
-We would be looking for the solutions of (55) and (60) in terms of elementary functions.
 Following the well known in mathematics principle, stating that the more general the problem is, the easier it is to solve it, let us change gears and instead of (60) consider the generalized damped Helmholtz-Duffing equation kogan4
@@
 VIII Back to the transmission line
-Now let us return to the particular cases of Eq.
-(49) (with the boundary conditions (39a)) considered in Section VI.
-For (60) using the results of the previous Section we claim that when the boundary conditions satisfy the relation
+Now let us return to the transmission line. For (60), using the results of the previous Section, we claim that when the parameters of the equation satisfy the relation
 $\frac{\tau_{R}^{2}}{T^{2}}=\frac{\gamma\left(Q_{1}+Q_{2}+2Q_{3}\right)^{2}}{2},$ (78)
@@
 $\chi=\sqrt{\frac{\beta(Q_{1}-Q_{2})}{6}}\cdot\frac{1}{T}.$ (83)
-Note that neither (78) goes to (81) nor (79) goes to (82) when $\gamma\to 0$.
-Small change in the equation leads to big change of the condition for the solution to be expressed in elementary functions.
+Note that (79) doesn’t goes to (82) when $\gamma\to 0$.
+More specifically, when $\gamma\to 0$, the solution (79) goes to the weak shock solution.
+On the other hand, the solution (82) corresponds to the case when both terms in (55) are of the same order of magnitude.
 We used previously the linear approximation for $V(Q)$ in the last term in the l.h.s.
@@
 of (49) in quadratic approximation (for cubic nonlinearity of $V(Q)$).
-The modification of Eqs.
-(78), (79) can be easily obtained from Eqs.
-(76), (77) for $n=1$.
+As the result, Eqs.
+(78) and (79) slightly change and Eq.
+(79) doesn’t change at all.
 VIII.1 The shocks vs. the kinks
@@
 It is reassuring to realize that the conditions for the integrability in elementary functions of Eqs.
-(55) and (60) describing weak shocks in half-nonlinear transmission line (i.e.
+(55) and (60) describing the shocks in half-nonlinear transmission line (i.e.
 Eqs.